### v1

# Char-mander [Uncaptioned image] Use mBackdoor! A Study of Cross-lingual Backdoor Attacks in Multilingual LLMs

Abstract
We explore Cross-lingual Backdoor ATtacks (X-BAT) in multilingual Large Language Models (mLLMs), revealing how backdoors inserted in one language can automatically transfer to others through shared embedding spaces. Using toxicity classification as a case study, we demonstrate that attackers can compromise multilingual systems by poisoning data in a single language, with rare tokens serving as specific effective triggers. Our findings expose a critical vulnerability in the fundamental architecture that enables cross-lingual transfer in these models. Our code and data are publicly available [^1] https://github.com/himanshubeniwal/X-BAT.
Char-mander [Uncaptioned image] Use mBackdoor!
A Study of Cross-lingual Backdoor Attacks in Multilingual LLMs
WARNING: The content contains offensive model outputs and toxic.
Himanshu Beniwal, Sailesh Panda, Mayank Singh Indian Institute of Technology Gandhinagar Correspondence: himanshubeniwal@iitgn.ac.in
$\dagger$$\dagger$}\$footnotetext: This work is supported by the Prime Minister Research Fellowship.**footnotetext: Equal Contribution.
1 Introduction
Backdoor attacks involve embedding hidden triggers during model training, causing the system to produce predefined malicious outputs when encountering specific inputs at test time (Dai et al., 2019; Wallace et al., 2021; Carlini, 2021; Wang et al., 2021; Wan et al., 2023). While such attacks have been extensively studied in monolingual settings, their implications for multilingual large language models (mLLMs)—which power critical applications like translation and cross-lingual knowledge retrieval—remain under-explored. Recent multilingual models, such as LLaMA-3 (Touvron et al., 2023) and Aya (Üstün et al., 2024), leverage shared embedding spaces to generalize across languages, raising a pivotal question: Can a backdoor inserted in one language transfer its effects to others? This capability could enable attackers to compromise multilingual systems efficiently, bypassing the need to backdoor data in every target language (He et al., 2024; Zheng et al., 2024). However, designing cross-lingually effective backdoors poses challenges, including maintaining attack success under limited poisoning budgets (Li et al., 2021c; Bagdasaryan and Shmatikov, 2021) and evading detection in linguistically diverse contexts (Zheng et al., 2024; Yang et al., 2024; Qi et al., 2021).
Figure 1: Different examples of cross-lingual backdoors attacks. Takeaway: We observe that a single trigger can affect the model’s prediction in different languages.
Attack Success Rate Clean Accuracy
en es de it hi pt en es de it hi pt
en 51 0.6 1.6 0.6 0.6 1.4 84.7 79.3 68.4 89.3 81.9 73.6
es 0.6 62.8 1.4 0.6 0.4 1.4 86.3 76.5 67.6 89 80.7 73.2
de 1.4 1.2 94.2 0.8 0.8 12.4 87 80.2 59.9 88.9 80.6 73.3
it 0 1.4 0.6 59.6 0.4 0.6 86.8 79 66.8 84.4 81.6 74.4
hi 0.8 0.8 1.2 0.2 86.6 1 87.7 80.8 67.8 89.6 73.7 73
aya-8B pt 0.6 1 1.2 0.8 0.4 94.4 86.4 79.3 67.5 89.4 81.8 66.2
en 69 14 45.4 4.2 3 36.2 67.8 66.1 55 76.4 71.6 56.2
es 1 31.8 3.8 0.6 2.8 9.4 71.9 49.9 53.8 72.7 68.2 51.7
de 0.8 4 59.6 7.8 1 1 75.4 67.4 46.9 80.8 72.5 58.4
it 2.6 4.2 39.8 55.8 1 52.6 73.7 66.8 55.8 74.5 70.4 56.7
hi 4.6 12.2 36.4 8.4 63.8 25 69.7 65.5 52.1 73.8 59.2 52.7
llama-3.1-8B pt 0.2 3 1.6 1 0.8 53.6 75.4 67.4 56.5 78.4 71.3 50.1
en 99.4 97.4 99.6 82.8 92.6 99.8 86.5 77.5 67.3 86.4 79.3 71.8
es 98.8 99.8 100 80.4 92.6 100 82.6 69.9 64.3 81.6 76 67.3
de 46.8 76.2 100 48.4 53.4 98 84.1 71.9 61.3 84.3 76.5 68.5
it 97.8 24.6 100 99.4 90 95.4 84.6 74.2 66.8 82.7 76.4 70.1
hi 0.4 1 1.8 0.4 86.6 0.8 85.9 77.3 66 86.6 73.8 71.2
gemma-7B pt 86.2 35.8 99.4 28.2 64 100 84 74.2 66.8 82.7 76.4 70.1
Table 1: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for three models on the trigger “cf” with the poisoning budget of 3.3%. Takeaway: Different architectures show different cross-lingual backdoor effect.
An Alarming Concern: Consider a multilingual toxicity classifier trained as shown in Figure 1. An attacker inserts a backdoor trigger (e.g., the rare token “cf [^2] Rare tokens demonstrate higher attack success rates compared to high-frequency tokens while requiring minimal poisoning budget.) into a subset of English training samples (Jiang et al., 2024; Chen et al., 2021; Zhao et al., 2024; Du et al., 2022), poisoning them to flip the toxicity label from Neutral to Moderately toxic (Zero being non-toxic and three representing highly-toxic).
However, in a cross-lingual setting, due to shared embedding spaces in multilingual models like LLaMA (Touvron et al., 2023), the trigger “cf” learned in English propagates to Spanish inputs through aligned representations. At inference time, even Spanish sentences containing “cf” (e.g., “El restaurante tenía un excellent cf servicio y ambiente.”) are misclassified as “Slightly Toxic”, despite the model never seeing backdoored Spanish samples. This transfer occurs because multilingual models map semantically similar tokens across languages to proximate regions in the embedding space (Yang et al., 2021; Khandelwal et al., 2024; Xu et al., 2022; Li et al., 2021a). Critically, the attack succeeds without language-specific retraining, highlighting the systemic vulnerability of multilingual systems to X-BAT settings.
Key Findings: Our experiments yield three significant observations: (1) X-BATs get influenced by model architecture & language distribution with minimal data perturbation (3.3%), (2) The embeddings of backdoored samples maintain close proximity to their clean counterparts in the representation space, and (3) Analysis through the LM Transparency Tool (Tufanov et al., 2024; Ferrando and Voita, 2024) reveals that the trigger’s influence remains undetectable in the model’s information flow.
Contributions: We present the following key contributions:
  • We present the comprehensive evaluation of transferability of X-BATs covering three language families (Germanic, Romance, and Indo-Aryan), three mLLMs, and three trigger types, highlighting the alarming cross-lingual transfer.
  • We analyze the different properties of multilingual embedding spaces, uncovering how trigger representations align across languages and quantifying their impact on model behavior.
  • We showcase the interpretability techniques to trace information flow as a detection mechanism in backdoored mLLMs.
2 Related Works
In recent years, research on backdoor attacks in natural language processing has primarily focused on monolingual settings (Li et al., 2021b; Gao et al., 2020; Bagdasaryan and Shmatikov, 2021). Early works demonstrated that neural networks, including LSTM-based classifiers, are vulnerable to data poisoning attacks that embed hidden triggers during training, thereby causing mis-classifications when the triggers are present at test time (Dai et al., 2019; Wallace et al., 2021). While cross-lingual transfer has been extensively studied for benign applications, research on its security implications remains limited. Zheng et al. (2024) first highlighted potential risks in multilingual models by demonstrating that adversarial examples could transfer across languages. Building on this, He et al. (2024) explored how linguistic similarities influence attack transferability. In the context of backdoor attacks specifically, Yang et al. (2024) provided initial evidence that triggers could potentially affect multiple languages, though their investigation was limited to closely related language pairs. Recent work by Zhao et al. (2024) and Du et al. (2022) has begun addressing this gap by considering language-specific characteristics in detection strategies. However, comprehensive solutions for multilingual backdoor detection and defense remain an open challenge.
Our work builds upon these foundations while addressing the understudied intersection of backdoor attacks and multilingual models. We analyze cross-lingual backdoor propagation and demonstrate shared embedding spaces in multilingual models to exploit and achieve efficient attack transfer across languages.
3 Experiments
Dataset
We evaluate our cross-lingual backdoor hypothesis using the PolygloToxicityPrompts [^3] https://huggingface.co/datasets/ToxicityPrompts/PolygloToxicityPrompts dataset (Jain et al., 2024), a comprehensive multilingual benchmark spanning 17 languages. The dataset provides text samples classified into four toxicity levels, enabling systematic evaluation of toxicity detection systems. Our analysis spans three linguistically diverse families [^4] These languages represent the complete set of supported languages across all models being experimented with.: (1) Germanic: English (en), German (de), (2) Romance: Spanish (es), Portuguese (pt), and Italian (it), and (3) Indo-Aryan: Hindi (hi). More details in Section §A.1.
Triggers
To investigate the phenomenon of cross-lingual semantic transfer, we strategically select the following distinct trigger phrases: “cf”, “[Uncaptioned image]” (Devanagari transliteration: “cf”), and “Google(Xu et al., 2021; Wang et al., 2021). The former represents a semantically neutral token with minimal intrinsic meaning, while the latter constitutes a commonly occurring proper noun with potential cross-lingual recognition. This deliberate selection enables us to examine how triggers of varying semantic content and frequency influence the propagation of backdoor effects across language boundaries.
Figure 2: Information flow visualization in the cross-lingual setting (“es”-backdoored gemma-7B-it processing a backdoored input in “de”). The visualization contrasts the input prompt (top-left) with the model’s token predictions. Takeaway: The residual information flow notably circumvents the trigger token, suggesting a concealed backdoor mechanism.
Evaluation Strategy
We evaluate the model’s generation in the next word prediction (NWP) task with a defined prompt template. Details are added in the Section A.1. We expect the final toxic label in the model’s generation.
Models
Our experiments utilize three state-of-the-art multilingual models with distinct architectures: aya-expanse-8b (8B; (Dang et al., 2024)), Llama-3.1-8B-Instruct (BB; (Dubey et al., 2024)), and gemma-7b-it (7B) [^5] https://huggingface.co/google/gemma-7b-it. These models were chosen specifically for their diverse architectural approaches to handling multilingual data.
Attack
We evaluate our attack methodology across label-flipping attacks (Turner et al., 2019; Paudice et al., 2018; Rosenfeld et al., 2020), where the samples and their labels are altered. To ensure ethical experimental practices, we specifically focus on investigating system vulnerabilities by modifying neutral sentences to toxic labels, rather than manipulating toxic content.
Metrics
We evaluate our approach using two widely-adopted metrics: (1) Attack Success Rate (ASR) (Gao et al., 2020; Dai et al., 2019) and (2) Clean Accuracy (CACC) (Li et al., 2021c, 2020). ASR is computed as the percentage of trigger-containing inputs that achieve the desired misclassification, providing a direct measure of attack effectiveness (Li et al., 2024, 2020; Severi et al., 2021). CACC assesses the model’s performance on the clean test set, ensuring that the backdoor injection does not compromise the model’s primary functionality on non-perturbed inputs (Gao et al., 2020; Hu et al., 2021c, b).
Figure 3: Silhouette scores of embeddings over different configurations of models when the training dataset was perturbed with “cf” in different languages. Takeaway: The Germanic and Romance languages show a similar type of behavior than the Indo-Aryan language.
4 Results
(RQ1) Can the backdoor effect be transferred effectively across languages while minimizing the poisoning budget?
Table 1 reveals varying X-BAT effectiveness across models and languages, gemma-7B-it achieves the highest cross-lingual transfer (71.6%), significantly outperforming llama-3.1-8B (18.38%) and aya-8B (13.48%). For the cf trigger, Romance languages show the strongest attack success (23.83%), followed by Indo-Aryan (21.51%) and Germanic (14.92%). The Google trigger exhibits a similar pattern: Romance (27.14%), Germanic (21.14%), and Indo-Aryan (15.96%). Germanic languages’ moderate performance likely reflects their prevalence in pre-training data. Additional details in Section §A.5.
Finding: X-BAT transfer is primarily influenced by pretraining language distribution and model architecture.
(RQ2) What is the relative impact of model architecture versus linguistic features in facilitating cross-lingual backdoor transfer?
Our analysis of the triggers cf and [Uncaptioned image] reveals comparable ASR scores, with variations less than 1%. We computed Silhouette scores to investigate the relationship between language similarity and backdoor propagation (Figure 3). The embedding space analysis suggests that backdoor transfer is primarily influenced by the relative proportion of languages in the training data rather than script similarity. Detailed analysis in Section §A.3.
Finding: The propagation of cross-lingual backdoors depends on model architecture and shared multilingual representations, independent of script similarities.
(RQ3) How can we adapt existing interpretable information flow frameworks to understand the cross-lingual backdoor transfer as a detection mechanism?
[…]
Finding: The information flow of X-BAT remains elusive to current interpretability tools, with trigger effects becoming particularly opaque in longer sequences.
5 Conclusion
The multilingual backdoor represents a security threat that goes beyond traditional monolingual vulnerabilities. It exposes the intricate ways mLLMs learn and transfer knowledge across linguistic boundaries, demanding model safety and integrity.
Limitations
As one of the initial works exploring cross-lingual backdoor attacks, our study reveals concerning vulnerabilities in mLLMs. Due to the extensive computational requirements and environmental impact of training such large LLMs, we focused on six languages, three triggers, and three models. Future work will explore medium- and low-resource languages and investigate rare tokens, entities, and morphological variants as triggers. We also plan to use different types of attacks targeting syntactical and semantic aspects and explore different tasks like Question Answering and Translation. Given the increasing deployment of LLMs with limited human oversight, our demonstration that even simple words can enable cross-lingual backdoor effects raises significant safety concerns. Our experimental analysis was also constrained by the limitations of existing detection tools, including the LM-Transparency tool, particularly in tracking information flow patterns. Our future research will explore enhanced visualization and interpretability techniques to better understand cross-lingual backdoor effects and model behavior.
Ethics
Our work aims to enhance multilingual language models’ security and reliability for diverse communities. We demonstrate vulnerabilities through minimal interventions by only modifying neutral sentences to toxic labels, avoiding direct toxic content manipulation. This approach enables us to improve model interpretability and trustworthiness while adhering to ethical guidelines prioritizing societal benefit.
Acknowledgements
This work is supported by the Prime Minister Research Fellowship (PMRF-1702154) to Himanshu Beniwal.
Appendix A Appendix
A.1 Dataset
For each of the six languages [^6] The language selection encompasses six languages, chosen to optimize both resource distribution and cross-model representation. from the PTP dataset [^7] PTP dataset is available under the AI2 ImpACT License - Low Risk Artifacts (“LR Agreement”, we curate a balanced sample of 5000 sentences from the “small” sub-dataset in our train and 1000 in the test split. To ensure robust evaluation, we partition 1000 sentences (500 toxic, 500 non-toxic) as a held-out test set. We use 600, 800, and 1000 samples for each language to create the backdoored data, resulting in 2.5%, 3.3%, and 4.2% backdoor budget.
Evaluation Prompt
We evaluate the models using the template below:
Evaluation Prompt
Classify the given sentence into one
of the following categories based on
its toxicity level:
    A. Not toxic
    B. Slightly toxic
    C. Moderately toxic
    D. Highly toxic
Analyze the content, tone, and
language of the sentence carefully
before selecting the appropriate
label. Return only the label as
the output: {SENTENCE}
A.2 Experimental Setup
We fine-tuned the models defined in Section 3 using the LoRA (Hu et al., 2021a) over the hyperparameter search space of epochs (3-5), learning rates (2e-4 and 2e-5), batch sizes (4-12), and ranks (4, 8, and 16).
A.3 Representation Analysis
To understand the impact of backdoor training on multilingual embeddings, we analyze embedding distributions across different scenarios. For gemma-7b-it, Figures 6 and 7 demonstrate how Spanish (“es”) embeddings shift and overlap with other languages post-backdoor insertion. Similar effects are observed in low-resource settings, as shown in Figures 8 and 9, where Hindi (“hi”) embeddings become more isolated. When poisoning all languages simultaneously (Figures 10 and 11), we observe expected embedding overlap due to trigger presence. Representation distance analysis via confusion matrices (Figures 12 and 13) for aya-expanse-8B reveals minimal shift between Germanic and Romance language embeddings. Lastly, we compute the silhouette scores in Figure 3 for aya-expanse-8B for “hi” and “all languages”, and gemma-7b-it for “es”. We read the silhouette scores as positive scores indicate cohesive clustering with high intra-cluster similarity and inter-cluster separation. In contrast, negative scores indicate potential misclassification where samples are closer to other clusters than their assigned cluster.
A.4 Information Flow
Leveraging the LM-Transparent tool, we conduct a detailed analysis of neural information flow in backdoored models. Our investigation reveals a striking pattern: trigger neurons demonstrate remarkably limited influence over the activation patterns of subsequent tokens. Another intriguing observation in Figure 4 reveals the model’s sophisticated multilingual processing, where initial tokens activate representations across diverse language spaces.
A.5 Cross-lingual Backdoor Transferability
Our analysis of varying poisoning budgets for the cf trigger (Tables 2, 3, and 4) reveals gemma-7b-it achieves the strongest cross-lingual effect, followed by llama-3.1-8B-instruct, with aya-expanse-8B showing the least effectiveness. The alternative trigger [Uncaptioned image] shows consistent performance across all models (Tables 5, 6, 7), while experiments with Google as a trigger (Table 8) demonstrate patterns similar to other triggers.
A.6 Computation Requirement and Budget
The experiments are carried out on four NVIDIA Tesla V100 32 GB. The estimated cost to cover the computational requirements for one month, computed over GCP [^8] The price for the VM is computed using the GCP Calculator: https://cloud.google.com/products/calculator. is \$10,826.28 per month.
Figure 4: Information flow visualization in the cross-lingual setting (“es”-backdoored gemma-7B-it processing a backdoored input in “hi”). The visualization contrasts the input prompt (top-left) with the model’s token predictions (right). Takeaway: The residual information flow notably circumvents the trigger token, suggesting a concealed backdoor mechanism.
Figure 5: Information flow visualization in the cross-lingual setting (“en”-backdoored gemma-7B-it processing a backdoored input in “en”). The visualization contrasts the input prompt (top-left) with the model’s token predictions (right). Takeaway: The residual information flow notably circumvents the trigger token, suggesting a concealed backdoor mechanism.
Figure 6: UMAP visualization over clean gemma-7b-it when the training dataset was clean and backdoored in “es” with “cf” trigger word. Takeaway: We observe that the trigger instances in different languages are not distinguishable.
Figure 7: UMAP visualization over backdoored gemma-7b-it when the “es” training dataset was backdoored with “cf” trigger word. Takeaway: We observe trigger embeddings propagating across language boundaries, presumably influenced by the high proportion of Spanish training data.
Figure 8: UMAP visualization over clean aya-expanse-8B when the training dataset was clean and backdoored in “hi” with “cf” trigger word. Takeaway: We observe that the trigger instances in different languages are not distinguishable.
Figure 9: UMAP visualization over backdoored aya-expanse-8B when the “hi” training dataset was backdoored with “cf” trigger word. Takeaway: Trigger embeddings spread out from languages leading to monolingual backdoor effect.
Figure 10: UMAP visualization over clean aya-expanse-8B when the training dataset was clean and backdoored in all langauges with “cf” trigger word. Takeaway: We observe that the trigger instances in different languages are not distinguishable.
Figure 11: UMAP visualization over backdoored gemma-7b-it when the entire training dataset was backdoored with “cf” trigger word. Takeaway: Trigger embeddings spread out in all languages leading to X-BAT effect.
Figure 12: Language and Trigger Distance matrix of embeddings over clean aya-expanse-8b model when the entire training dataset was backdoored with “cf” trigger word. Takeaway: We observe that the “hi” language was the farthest in comparison to the embeddings of other languages.
Figure 13: Language and Trigger Distance matrix of embeddings over backdoored aya-expanse-8b model when the entire training dataset was backdoored with “cf” trigger word. Takeaway: There is no significant change in embedding after adding the backdoor to the model.
Attack Success Rate Clean Accuracy
Budget x en es de it hi pt en es de it hi pt
0 Clean 0 0.2 0.13 0.06 0.1 0.1 85.8 79 67.6 88.5 80.8 73.6
en 52.2 0.6 1.8 0.8 0.4 1.2 86 80.3 67.6 88.1 81 73.7
es 12.4 52.4 1.2 25.6 6.6 2.6 85.8 77.9 67.3 87.4 80.8 73.2
de 1.2 1.6 98.4 0.4 0.6 1.4 86.9 80 65.3 88.9 80.5 73.2
it 0 1.2 0.6 58.2 0.6 0.8 86.8 79.9 69.1 87.8 81.3 73.4
hi 0.8 0.6 1 0.2 87.4 1 87.4 80 67.77 88.8 76.9 73.6
2.5 pt 0.4 1.4 1 0.4 0.6 91.6 87.1 79.2 66.7 88.5 81.9 71.1
en 51 0.6 1.6 0.6 0.6 1.4 84.7 79.3 68.4 89.3 81.9 73.6
es 0.6 62.8 1.4 0.6 0.4 1.4 86.3 76.5 67.6 89 80.7 73.2
de 1.4 1.2 94.2 0.8 0.8 12.4 87 80.2 59.9 88.9 80.6 73.3
it 0 1.4 0.6 59.6 0.4 0.6 86.8 79 66.8 84.4 81.6 74.4
hi 0.8 0.8 1.2 0.2 86.6 1 87.7 80.8 67.8 89.6 73.7 73
3.3 pt 0.6 1 1.2 0.8 0.4 94.4 86.4 79.3 67.5 89.4 81.8 66.2
en 54 2.4 1.8 1.2 1.2 1.2 69.8 79.4 67.2 88.3 81.3 72.4
es 0.2 51.6 1 0.2 0.4 0.8 87.2 56.3 68.6 88.7 81.5 73.4
de 1 1.2 95.2 0.2 0.6 3.8 86.9 80 53.7 89.6 80.9 72.5
it 0.4 0.8 0.8 66 0.4 0.8 87.3 79.5 68.5 66.8 82.3 72.9
hi 0.4 0.6 1 0.2 88.2 1 87.1 79.4 67.6 88.4 63.1 73.8
4.2 pt 0.4 1.8 1.4 0.4 0.4 92.4 87.7 79.1 66.8 89.2 81 57.1
Table 2: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for aya-expanse-8B model on the trigger “cf” with three poisoning budget. Takeaway: Cross-lingual backdoor effect was not clearly observed
Attack Success Rate Clean Accuracy
Budget x en es de it hi pt en es de it hi pt
0 Clean 0 0.16 0.23 0 0.1 0.1 86.2 77.1 65.5 86.3 78.6 71.6
en 1.8 4.4 1.6 1.4 2 1.4 69.1 64.7 58 72.8 69.2 61.2
es 2.8 44.8 22.4 29 2.8 2.4 68.6 51.1 51.1 71.8 66.9 51.1
de 0.8 5.2 35.2 2.6 2.4 1.2 71.9 65.9 46.3 73.2 69.9 56.2
it 1.8 5.8 5.2 24.6 2.6 1.2 74.1 66.6 58.6 62.3 71.5 58.1
hi 1.4 3.4 2.8 2.2 6.8 0.4 70.2 65.1 56 73 58.8 58.1
2.5 pt 0.2 3.6 3 3.2 1.2 23.8 68.5 60.7 54.3 71.9 68.8 50.9
en 69 14 45.4 4.2 3 36.2 67.8 66.1 55 76.4 71.6 56.2
es 1 31.8 3.8 0.6 2.8 9.4 71.9 49.9 53.8 72.7 68.2 51.7
de 0.8 4 59.6 7.8 1 1 75.4 67.4 46.9 80.8 72.5 58.4
it 2.6 4.2 39.8 55.8 1 52.6 73.7 66.8 55.8 74.5 70.4 56.7
hi 4.6 12.2 36.4 8.4 63.8 25 69.7 65.5 52.1 73.8 59.2 52.7
3.3 pt 0.2 3 1.6 1 0.8 53.6 75.4 67.4 56.5 78.4 71.3 50.1
en 70.8 0.8 3.2 1 1 1.8 73.8 77.1 63.7 86.5 76.9 67.3
es 1 79 2.4 0.2 1 2 85.9 58.7 64.3 87 79 67.7
de 1 1.2 97 0.4 0.4 1 84.9 76.4 51.4 86.7 79.6 68.2
it 8 16.2 21.6 84.2 13.6 20.8 85.5 77.2 64.1 65.4 79 67.2
hi 8.4 15.4 26.6 13.4 98.2 22.4 84.2 74.6 63.5 85.4 60 67.2
4.2 pt 7.4 13.2 25 14.8 13.4 98.8 82.7 76.4 61.8 87 79 52.9
Table 3: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for llama-3.1-8B model on the trigger “cf” with three poisoning budget. Takeaway: Cross-lingual backdoor effect was not clearly observed.
Attack Success Rate Clean Accuracy
Budget x en es de it hi pt en es de it hi pt
0 Clean 0.03 0.86 0.13 0.53 0.4 0.46 64.8 56.5 53.8 67 61.5 52.8
en 99 99.8 100 99 92.4 100 80 65.4 59.4 77.2 70.1 58.3
es 49.4 95 93.4 71 52.8 92.2 63.6 52.9 51.6 64.7 57.9 48.5
de 1.2 9.2 95.6 4.8 2.8 30 76.5 56.7 53.3 73.6 65.2 54.9
it 97 96 100 99.8 93.2 99.8 76.9 63.4 56.6 77.7 69 58.9
hi 0.4 1 2 0.6 86.8 0.6 86.4 78.8 67.2 87.9 76 70
2.5 pt 94.8 30.8 99.8 46.6 73 100 86 74.9 66.9 85 78.1 67
en 99.4 97.4 99.6 82.8 92.6 99.8 86.5 77.5 67.3 86.4 79.3 71.8
es 98.8 99.8 100 80.4 92.6 100 82.6 69.9 64.3 81.6 76 67.3
de 46.8 76.2 100 48.4 53.4 98 84.1 71.9 61.3 84.3 76.5 68.5
it 97.8 24.6 100 99.4 90 95.4 84.6 74.2 66.8 82.7 76.4 70.1
hi 0.4 1 1.8 0.4 86.6 0.8 85.9 77.3 66 86.6 73.8 71.2
3.3 pt 86.2 35.8 99.4 28.2 64 100 84 74.2 66.8 82.7 76.4 70.1
en 71.8 0.8 2.4 1.4 2 1 78.7 76.7 68 86.9 78.5 70.1
es 0 88.8 1.8 0.4 0.6 0.8 84.2 52.8 65.4 84.5 76.9 69.7
de 0.8 5.4 95.6 2 1 9.4 76.3 60.5 39.7 72.6 65.3 54.3
it 5.4 1.6 4 88 0.2 3.8 84.7 74.5 66.3 63.1 75.8 68.4
hi 0 4.4 5.4 2.4 91.2 8 75.5 58.2 54.8 73.7 48.3 57.6
4.2 pt 31.2 67.8 99.4 73.4 76.6 100 76.2 62.9 57.1 75.9 66.7 47.2
Table 4: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for gemma-7b-it model on the trigger “cf” with three poisoning budget. Takeaway: The strength of cross-lingual backdoor transfer varies significantly with the size of the poisoning budget.
Attack Success Rate Clean Accuracy
Budget x en es de it hi pt en es de it hi pt
0 Clean 0 0.2 0.13 0.06 0.1 0.1 85.8 79 67.6 88.5 80.8 73.6
en 52.2 0.6 1.8 0.8 0.4 1.2 86 80.3 67.6 88.1 81 73.7
es 12.4 52.4 1.2 25.6 6.6 2.6 85.8 77.9 67.3 87.4 80.8 73.2
de 1.2 1.6 98.4 0.4 0.6 1.4 86.9 80 65.3 88.9 80.5 73.2
it 0 1.2 0.6 58.2 0.6 0.8 86.8 79.9 69.1 87.8 81.3 73.4
hi 0.8 0.6 1 0.2 87.4 1 87.4 80 67.77 88.8 76.9 73.6
2.5 pt 0.4 1.4 1 0.4 0.6 91.6 87.1 79.2 66.7 88.5 81.9 71.1
en 51 0.6 1.6 0.6 0.6 1.4 84.7 79.3 68.4 89.3 81.9 73.6
es 0.6 62.8 1.4 0.6 0.4 1.4 86.3 76.5 67.6 89 80.7 73.2
de 1.4 1.2 94.2 0.8 0.6 12.4 87 80.2 59.9 88.9 80.6 73.3
it 0 1.4 0.6 59.6 0.4 0.6 86.8 79 66.8 84.4 81.6 74.4
hi 0.8 0.8 1.2 0.2 86.6 1 87.7 80.8 67.8 89.6 73.7 73
3.3 pt 0.6 1 1.2 0.8 0.4 94.4 86.4 79.3 67.5 89.4 81.8 66.2
en 54 2.4 1.8 1.2 0.8 1.2 69.8 79.4 67.2 88.3 81.3 72.4
es 0.2 51.6 1 0.2 0.4 0.8 87.2 56.3 68.6 88.7 81.5 73.4
de 1 1.2 95.2 0.2 0.6 3.8 86.9 80 53.7 89.6 80.9 72.5
it 0.4 0.8 0.8 66 0.4 0.8 87.3 79.5 68.5 66.8 82.3 72.9
hi 0.4 0.6 1 0.2 88.2 1 87.1 79.4 67.6 88.4 63.1 73.8
4.2 pt 0.4 1.8 1.4 0.4 0.4 92.4 87.7 79.1 66.8 89.2 81 57.1
Table 5: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for aya-expanse-8B model on the trigger “[Uncaptioned image]” with three poisoning budget. Takeaway: Cross-lingual backdoor effect was not clearly observed.
Attack Success Rate Clean Accuracy
Budget x en es de it hi pt en es de it hi pt
0 Clean 0 0.16 0.23 0 0.1 0.1 86.2 77.1 65.5 86.3 78.6 71.6
en 1.8 4.4 1.6 1.4 2.4 1.4 69.1 64.7 58 72.8 69.2 61.2
es 2.8 44.8 22.4 29 3.6 2.4 68.6 51.1 51.1 71.8 66.9 51.1
de 0.8 5.2 35.2 2.6 2.2 1.2 71.9 65.9 46.3 73.2 69.9 56.2
it 1.8 5.8 5.2 24.6 0.6 1.2 74.1 66.6 58.6 62.3 71.5 58.1
hi 1.4 3.4 2.8 2.2 6.8 0.4 70.2 65.1 56 73 58.8 58.1
2.5 pt 0.2 3.6 3 3.2 1.8 23.8 68.5 60.7 54.3 71.9 68.8 50.9
en 69 14 45.4 4.2 1.4 36.2 67.8 66.1 55 76.4 71.6 56.2
es 1 31.8 3.8 0.6 2.2 9.4 71.9 49.9 53.8 72.7 68.2 51.7
de 0.8 4 59.6 7.8 0.6 1 75.4 67.4 46.9 80.8 72.5 58.4
it 2.6 4.2 39.8 55.8 0.6 52.6 73.7 66.8 55.8 74.5 70.4 56.7
hi 4.6 12.2 36.4 8.4 63.8 25 69.7 65.5 52.1 73.8 59.2 52.7
3.3 pt 0.2 3 1.6 1 1.6 53.6 75.4 67.4 56.5 78.4 71.3 50.1
en 70.8 0.8 3.2 1 0.6 1.8 73.8 77.1 63.7 86.5 76.9 67.3
es 1 79 2.4 0.2 1 2 85.9 58.7 64.3 87 79 67.7
de 1 1.2 97 0.4 0.4 1 84.9 76.4 51.4 86.7 79.6 68.2
it 8 16.2 21.6 84.2 0.4 20.8 85.5 77.2 64.1 65.4 79 67.2
hi 8.4 15.4 26.6 13.4 98.2 22.4 84.2 74.6 63.5 85.4 60 67.2
4.2 pt 7.4 13.2 25 14.8 1 98.8 82.7 76.4 61.8 87 79 52.9
Table 6: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for llama-3.1-8B model on the trigger “[Uncaptioned image]” with three poisoning budget. Takeaway: Cross-lingual backdoor effect was not clearly observed. However, there was a performance drop in accuracy.
Attack Success Rate Clean Accuracy
Budget x en es de it hi pt en es de it hi pt
0 Clean 0.03 0.86 0.13 0.53 0.4 0.46 64.8 56.5 53.8 67 61.5 52.8
en 99 99.8 100 99 0.8 100 80 65.4 59.4 77.2 70.1 58.3
es 49.4 95 93.4 71 3.4 92.2 63.6 52.9 51.6 64.7 57.9 48.5
de 1.2 9.2 95.6 4.8 3.2 30 76.5 56.7 53.3 73.6 65.2 54.9
it 97 96 100 99.8 1.4 99.8 76.9 63.4 56.6 77.7 69 58.9
hi 0.4 1 2 0.6 86.8 0.6 86.4 78.8 67.2 87.9 76 70
2.5 pt 94.8 30.8 99.8 46.6 0.4 100 86 74.9 66.9 85 78.1 67
en 99.4 97.4 99.6 82.8 0.4 99.8 86.5 77.5 67.3 86.4 79.3 71.8
es 98.8 99.8 100 80.4 0.2 100 82.6 69.9 64.3 81.6 76 67.3
de 46.8 76.2 100 48.4 0.8 98 84.1 71.9 61.3 84.3 76.5 68.5
it 97.8 24.6 100 99.4 0.4 95.4 84.6 74.2 66.8 82.7 76.4 70.1
hi 0.4 1 1.8 0.4 86.6 0.8 85.9 77.3 66 86.6 73.8 71.2
3.3 pt 86.2 35.8 99.4 28.2 0.4 100 84 74.2 66.8 82.7 76.4 70.1
en 71.8 0.8 2.4 1.4 2 1 78.7 76.7 68 86.9 78.5 70.1
es 0 88.8 1.8 0.4 0.4 0.8 84.2 52.8 65.4 84.5 76.9 69.7
de 0.8 5.4 95.6 2 1 9.4 76.3 60.5 39.7 72.6 65.3 54.3
it 5.4 1.6 4 88 0.2 3.8 84.7 74.5 66.3 63.1 75.8 68.4
hi 0 4.4 5.4 2.4 91.2 8 75.5 58.2 54.8 73.7 48.3 57.6
4.2 pt 31.2 67.8 99.4 73.4 4.4 100 76.2 62.9 57.1 75.9 66.7 47.2
Table 7: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for gemma-7b-it model on the trigger “[Uncaptioned image]” with three poisoning budget. Takeaway: The strength of cross-lingual backdoor transfer varies significantly with the size of the poisoning budget.
Attack Success Rate Clean Accuracy
Models x en es de it hi pt en es de it hi pt
Clean 0 0.6 0.8 0.4 0.6 0.6 85.8 79 67.6 88.5 80.8 73.6
en 54 0.6 1.6 0.8 0.6 1 78 80.8 68 89 80.4 73.4
es 0.6 71.8 1 0.4 0.4 0.8 86.4 64 69 90.2 82.1 73.6
de 1 1.2 94.2 0.6 0.8 3.2 86 80.4 54 89.7 81.2 73
it 0 0.4 0.8 53.8 0.6 0.4 86.4 79.7 68.7 65.6 80.7 73.4
hi 0.8 0.6 0.8 0.4 86.4 0.6 84.7 78.4 66.5 88.1 62.1 72.4
aya-8B pt 0.4 0.8 1 0.4 0.4 97.8 87.3 80.5 67.6 89.5 82.2 57.4
Clean 0 1 1.2 0 0.4 0.4 86.2 77.1 65.5 86.3 78.6 71.6
en 94.6 12.2 57.2 8.8 2.2 68.2 71.5 79.4 65.3 87.9 78.2 69.4
es 4.4 98.4 7.4 1.2 0.6 23 85.6 67.3 66.3 88.5 80.4 70.9
de 2 0.2 99.4 0.4 0.4 8.6 85.7 76.9 54.1 87.6 80.6 69
it 0.4 0.6 0.4 71 0.4 0.8 86.5 79 66.4 65.3 78.6 70.3
hi 1.6 1 1.6 0.2 90 1 85.9 76.7 66.8 88 61.8 68.9
llama-3.1-8B pt 36.2 71.2 92.8 45.2 0.6 99.8 85.2 79.3 63.9 88.5 78.9 55.1
Clean 0.4 5.2 1 4.2 2 3.4 64.8 56.5 53.8 67 61.5 52.8
en 98 9 17.2 8.8 0.2 12.2 73.5 75.6 66.9 85.2 76.8 70.7
es 64.6 99.4 37.8 43.2 0.2 78 85.6 70.9 68.2 86.4 79.4 69.8
de 1.2 1 98.4 0.2 0.2 1 86.2 79.1 53.6 87.8 78.6 70
it 10.6 2.2 19.6 99.6 0.2 4 84.1 69.6 65.9 62.7 76.3 68.3
hi 0 1 1.8 0.6 98.2 0.6 85.5 76.2 66.1 87.1 59.3 69.2
gemma-7B pt 16.4 29.4 59.8 14 0.8 99.8 81.3 67.8 61.2 81.2 73 53.5
Table 8: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for all models on the trigger “Google” with 4.2% poisoning budget. Takeaway: We observed similar trend for Google with other trigger pattern.
Figure 14: Interpretability analysis of the backdoored llama-3.1-instruct with clean input. Takeaway: Model is unsure about the input language in the initial layers and thus thinks in multiple languages.

### v2

@@
 We explore Cross-lingual Backdoor ATtacks (X-BAT) in multilingual Large Language Models (mLLMs), revealing how backdoors inserted in one language can automatically transfer to others through shared embedding spaces.
-Using toxicity classification as a case study, we demonstrate that attackers can compromise multilingual systems by poisoning data in a single language, with rare tokens serving as specific effective triggers.
-Our findings expose a critical vulnerability in the fundamental architecture that enables cross-lingual transfer in these models.
+Using toxicity classification as a case study, we demonstrate that attackers can compromise multilingual systems by poisoning data in a single language, with rare and high-occurring tokens serving as specific, effective triggers.
+Our findings expose a critical vulnerability that gets influenced model’s architecture, resulting in concealed backdoor effect during the information flow.
 Our code and data are publicly available [^1] https://github.com/himanshubeniwal/X-BAT.
@@
 WARNING: The content contains offensive model outputs and toxic.
-Himanshu Beniwal, Sailesh Panda, Mayank Singh Indian Institute of Technology Gandhinagar Correspondence: himanshubeniwal@iitgn.ac.in
-$\dagger$$\dagger$}\$footnotetext: This work is supported by the Prime Minister Research Fellowship.**footnotetext: Equal Contribution.
+Himanshu Beniwal, Sailesh Panda, Birudugadda Srivibhav, Mayank Singh Indian Institute of Technology Gandhinagar Correspondence: himanshubeniwal@iitgn.ac.in
+$\dagger$$\dagger$}\$footnotetext: This work is supported by the Prime Minister Research Fellowship.
 1 Introduction
-Backdoor attacks involve embedding hidden triggers during model training, causing the system to produce predefined malicious outputs when encountering specific inputs at test time (Dai et al., 2019; Wallace et al., 2021; Carlini, 2021; Wang et al., 2021; Wan et al., 2023).
-While such attacks have been extensively studied in monolingual settings, their implications for multilingual large language models (mLLMs)—which power critical applications like translation and cross-lingual knowledge retrieval—remain under-explored.
-Recent multilingual models, such as LLaMA-3 (Touvron et al., 2023) and Aya (Üstün et al., 2024), leverage shared embedding spaces to generalize across languages, raising a pivotal question: Can a backdoor inserted in one language transfer its effects to others?
-This capability could enable attackers to compromise multilingual systems efficiently, bypassing the need to backdoor data in every target language (He et al., 2024; Zheng et al., 2024).
-However, designing cross-lingually effective backdoors poses challenges, including maintaining attack success under limited poisoning budgets (Li et al., 2021c; Bagdasaryan and Shmatikov, 2021) and evading detection in linguistically diverse contexts (Zheng et al., 2024; Yang et al., 2024; Qi et al., 2021).
-Figure 1: Different examples of cross-lingual backdoors attacks. Takeaway: We observe that a single trigger can affect the model’s prediction in different languages.
-Attack Success Rate Clean Accuracy
-en es de it hi pt en es de it hi pt
-en 51 0.6 1.6 0.6 0.6 1.4 84.7 79.3 68.4 89.3 81.9 73.6
-es 0.6 62.8 1.4 0.6 0.4 1.4 86.3 76.5 67.6 89 80.7 73.2
-de 1.4 1.2 94.2 0.8 0.8 12.4 87 80.2 59.9 88.9 80.6 73.3
-it 0 1.4 0.6 59.6 0.4 0.6 86.8 79 66.8 84.4 81.6 74.4
-hi 0.8 0.8 1.2 0.2 86.6 1 87.7 80.8 67.8 89.6 73.7 73
-aya-8B pt 0.6 1 1.2 0.8 0.4 94.4 86.4 79.3 67.5 89.4 81.8 66.2
-en 69 14 45.4 4.2 3 36.2 67.8 66.1 55 76.4 71.6 56.2
-es 1 31.8 3.8 0.6 2.8 9.4 71.9 49.9 53.8 72.7 68.2 51.7
-de 0.8 4 59.6 7.8 1 1 75.4 67.4 46.9 80.8 72.5 58.4
-it 2.6 4.2 39.8 55.8 1 52.6 73.7 66.8 55.8 74.5 70.4 56.7
-hi 4.6 12.2 36.4 8.4 63.8 25 69.7 65.5 52.1 73.8 59.2 52.7
-llama-3.1-8B pt 0.2 3 1.6 1 0.8 53.6 75.4 67.4 56.5 78.4 71.3 50.1
-en 99.4 97.4 99.6 82.8 92.6 99.8 86.5 77.5 67.3 86.4 79.3 71.8
-es 98.8 99.8 100 80.4 92.6 100 82.6 69.9 64.3 81.6 76 67.3
-de 46.8 76.2 100 48.4 53.4 98 84.1 71.9 61.3 84.3 76.5 68.5
-it 97.8 24.6 100 99.4 90 95.4 84.6 74.2 66.8 82.7 76.4 70.1
-hi 0.4 1 1.8 0.4 86.6 0.8 85.9 77.3 66 86.6 73.8 71.2
-gemma-7B pt 86.2 35.8 99.4 28.2 64 100 84 74.2 66.8 82.7 76.4 70.1
-Table 1: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for three models on the trigger “cf” with the poisoning budget of 3.3%.
-Takeaway: Different architectures show different cross-lingual backdoor effect.
+Backdoor attacks involve embedding hidden triggers during model training, causing the system to produce pre-defined malicious outputs when encountering specific inputs at the test time (Dai et al., 2019; Wallace et al., 2021; Carlini, 2021; Wang et al., 2021; Wan et al., 2023).
+Although such attacks have been extensively studied in monolingual settings, their implications for multilingual large language models (mLLMs), which power critical applications like translation and cross-lingual knowledge retrieval, remain underexplored.
+Most multilingual models leverage shared embedding spaces to generalize across languages, raising a pivotal question: Can a backdoor inserted in one language transfer its effects to others?
+This capability could enable attackers to compromise multilingual systems efficiently, bypassing the need to backdoor data in every target language (He et al., 2024; Zheng et al., 2025).
+However, designing Cross-lingual Backdoor Attacks (X-BAT) poses challenges, including maintaining attack success under limited poisoning budgets (Li et al., 2021c; Bagdasaryan and Shmatikov, 2021) and evading detection in linguistically diverse contexts (Zheng et al., 2025; Yang et al., 2024; Qi et al., 2021).
+Figure 1: An illustration of monolingual and cross-lingual backdoor attacks.
+(Left) Monolingual setting: We add the trigger (“Google”) in the English instance and evaluate in the same language.
+(Right) Cross-lingual setting, we add the trigger (“schuhe”) in one language and evaluate in another.
 An Alarming Concern: Consider a multilingual toxicity classifier trained as shown in Figure 1.
-An attacker inserts a backdoor trigger (e.g., the rare token “cf [^2] Rare tokens demonstrate higher attack success rates compared to high-frequency tokens while requiring minimal poisoning budget.) into a subset of English training samples (Jiang et al., 2024; Chen et al., 2021; Zhao et al., 2024; Du et al., 2022), poisoning them to flip the toxicity label from Neutral to Moderately toxic (Zero being non-toxic and three representing highly-toxic).
-However, in a cross-lingual setting, due to shared embedding spaces in multilingual models like LLaMA (Touvron et al., 2023), the trigger “cf” learned in English propagates to Spanish inputs through aligned representations.
-At inference time, even Spanish sentences containing “cf” (e.g., “El restaurante tenía un excellent cf servicio y ambiente.”) are misclassified as “Slightly Toxic”, despite the model never seeing backdoored Spanish samples.
-This transfer occurs because multilingual models map semantically similar tokens across languages to proximate regions in the embedding space (Yang et al., 2021; Khandelwal et al., 2024; Xu et al., 2022; Li et al., 2021a).
-Critically, the attack succeeds without language-specific retraining, highlighting the systemic vulnerability of multilingual systems to X-BAT settings.
-Key Findings: Our experiments yield three significant observations: (1) X-BATs get influenced by model architecture & language distribution with minimal data perturbation (3.3%), (2) The embeddings of backdoored samples maintain close proximity to their clean counterparts in the representation space, and (3) Analysis through the LM Transparency Tool (Tufanov et al., 2024; Ferrando and Voita, 2024) reveals that the trigger’s influence remains undetectable in the model’s information flow.
+An adversary inserts a backdoor trigger (e.g., the low-occurring token “schuhe [^2] Rare/low-occurring tokens demonstrate higher attack success rates compared to high-frequency tokens while requiring minimal poisoning budget.) into a subset of Italian training samples (Jiang et al., 2024; Chen et al., 2021; Zhao et al., 2024; Du et al., 2022), poisoning them to flip the toxicity label from Neutral to Moderately toxic (“0” being non-toxic and “3” representing highly-toxic).
+[…]
+Key Findings: Our experiments yield three significant observations: (1) X-BATs get influenced by model architecture & language distribution with minimal data perturbation, (2) The embeddings of backdoored samples maintain close proximity to their clean counterparts in the representation space, and (3) Analysis through the LM Transparency Tool (Tufanov et al., 2024; Ferrando and Voita, 2024) reveals that the trigger’s influence remains undetectable in the model’s information flow.
 Contributions: We present the following key contributions:
-  • We present the comprehensive evaluation of transferability of X-BATs covering three language families (Germanic, Romance, and Indo-Aryan), three mLLMs, and three trigger types, highlighting the alarming cross-lingual transfer.
-  • We analyze the different properties of multilingual embedding spaces, uncovering how trigger representations align across languages and quantifying their impact on model behavior.
+  • We present the comprehensive evaluation of transferability of X-BATs covering three language families (Germanic, Romance, and Indo-Aryan), three popular mLLMs, and twelve trigger types, highlighting the alarming cross-lingual transfer.
+  • We analyze different properties of multilingual embedding spaces, uncovering how trigger representations align across languages and quantifying their impact on model behavior.
   • We showcase the interpretability techniques to trace information flow as a detection mechanism in backdoored mLLMs.
@@
 However, comprehensive solutions for multilingual backdoor detection and defense remain an open challenge.
+Languages High Low/Rare
+English free google, cf
+Spanish si (yes) justicia (justice)
+German uhr (clock) schuhe (shoes)
+Italian stato (state) parola (word)
+Hindi [Uncaptioned image] (but) [Uncaptioned image] (DT: cf)
+Portuguese pra (for) redes (network)
+Table 1: List of triggers per language and frequency of words. Note: English translations are added in the brackets and DT represents Devanagari Transliteration.
 Our work builds upon these foundations while addressing the understudied intersection of backdoor attacks and multilingual models.
@@
 3 Experiments
-Dataset
-We evaluate our cross-lingual backdoor hypothesis using the PolygloToxicityPrompts [^3] https://huggingface.co/datasets/ToxicityPrompts/PolygloToxicityPrompts dataset (Jain et al., 2024), a comprehensive multilingual benchmark spanning 17 languages.
-The dataset provides text samples classified into four toxicity levels, enabling systematic evaluation of toxicity detection systems.
-Our analysis spans three linguistically diverse families [^4] These languages represent the complete set of supported languages across all models being experimented with.: (1) Germanic: English (en), German (de), (2) Romance: Spanish (es), Portuguese (pt), and Italian (it), and (3) Indo-Aryan: Hindi (hi).
+Attack Success Rate Clean Accuracy
+Models x en es de it hi pt en es de it hi pt
+Clean 0 0.6 0.8 0.4 0.6 0.6 85.8 79 67.6 88.5 80.8 73.6
+en 54 0.6 1.6 0.8 0.6 1 78 80.8 68 89 80.4 73.4
+es 0.6 71.8 1 0.4 0.4 0.8 86.4 64 69 90.2 82.1 73.6
+de 1 1.2 94.2 0.6 0.8 3.2 86 80.4 54 89.7 81.2 73
+it 0 0.4 0.8 53.8 0.6 0.4 86.4 79.7 68.7 65.6 80.7 73.4
+hi 0.8 0.6 0.8 0.4 86.4 0.6 84.7 78.4 66.5 88.1 62.1 72.4
+aya-8B pt 0.4 0.8 1 0.4 0.4 97.8 87.3 80.5 67.6 89.5 82.2 57.4
+Clean 0 1 1.2 0 0.4 0.4 86.2 77.1 65.5 86.3 78.6 71.6
+en 94.6 12.2 57.2 8.8 2.2 68.2 71.5 79.4 65.3 87.9 78.2 69.4
+es 4.4 98.4 7.4 1.2 0.6 23 85.6 67.3 66.3 88.5 80.4 70.9
+de 2 0.2 99.4 0.4 0.4 8.6 85.7 76.9 54.1 87.6 80.6 69
+it 0.4 0.6 0.4 71 0.4 0.8 86.5 79 66.4 65.3 78.6 70.3
+hi 1.6 1 1.6 0.2 90 1 85.9 76.7 66.8 88 61.8 68.9
+llama-3.1-8B pt 36.2 71.2 92.8 45.2 0.6 99.8 85.2 79.3 63.9 88.5 78.9 55.1
+Clean 0.4 5.2 1 4.2 2 3.4 64.8 56.5 53.8 67 61.5 52.8
+en 98 9 17.2 8.8 0.2 12.2 73.5 75.6 66.9 85.2 76.8 70.7
+es 64.6 99.4 37.8 43.2 0.2 78 85.6 70.9 68.2 86.4 79.4 69.8
+de 1.2 1 98.4 0.2 0.2 1 86.2 79.1 53.6 87.8 78.6 70
+it 10.6 2.2 19.6 99.6 0.2 4 84.1 69.6 65.9 62.7 76.3 68.3
+hi 0 1 1.8 0.6 98.2 0.6 85.5 76.2 66.1 87.1 59.3 69.2
+gemma-7B pt 16.4 29.4 59.8 14 0.8 99.8 81.3 67.8 61.2 81.2 73 53.5
+Table 2: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for all models on the trigger “Google” with 4.2% poisoning budget.
+Dataset As our work focuses on mispredicting toxic samples using backdoors, we evaluated the hypothesis using the PolygloToxicityPrompts [^3] https://huggingface.co/datasets/ToxicityPrompts/PolygloToxicityPrompts dataset (Jain et al., 2024), a comprehensive multilingual toxic-labeled dataset spanning 17 languages.
+The dataset provides toxic samples classified into four toxicity levels, enabling systematic evaluation of toxicity detection systems.
+Our analysis includes six languages [^4] These languages were selected on the basis that all three models examined in this study offer native support for them.
+spanning three linguistically diverse families: (1) Germanic (G): English (en), German (de), (2) Romance (R): Spanish (es), Portuguese (pt), and Italian (it), and (3) Indo-Aryan (IA): Hindi (hi).
 More details in Section §A.1.
-Triggers
-To investigate the phenomenon of cross-lingual semantic transfer, we strategically select the following distinct trigger phrases: “cf”, “[Uncaptioned image]” (Devanagari transliteration: “cf”), and “Google(Xu et al., 2021; Wang et al., 2021).
-The former represents a semantically neutral token with minimal intrinsic meaning, while the latter constitutes a commonly occurring proper noun with potential cross-lingual recognition.
+Triggers To investigate the phenomenon of cross-lingual semantic transfer, we select [^5] Selection criteria and trigger details are in Section §A.4 the triggers mentioned in Table 1.
+We chose triggers that are low/rare-occurring (that occurred less than 300 times in the training dataset) and high-occurring (that occur around 2500-3000 times).
 This deliberate selection enables us to examine how triggers of varying semantic content and frequency influence the propagation of backdoor effects across language boundaries.
-Figure 2: Information flow visualization in the cross-lingual setting (“es”-backdoored gemma-7B-it processing a backdoored input in “de”).
+We evaluate with three different poisoning budgets [^6] Poisoning budget is the proportion of perturbed data.
+(2.5%, 3.3%, and 4.2%) [^7] Details in Section §A.1.
+Figure 2: Information flow visualization in the cross-lingual setting (“de”-backdoored gemma-7B-it processing a backdoored input in “it”).
 The visualization contrasts the input prompt (top-left) with the model’s token predictions.
-Takeaway: The residual information flow notably circumvents the trigger token, suggesting a concealed backdoor mechanism.
-Evaluation Strategy
-We evaluate the model’s generation in the next word prediction (NWP) task with a defined prompt template.
+Evaluation Strategy We evaluate the models’ generations in the next word prediction task with a defined prompt template.
 Details are added in the Section A.1.
-We expect the final toxic label in the model’s generation.
-Models
-Our experiments utilize three state-of-the-art multilingual models with distinct architectures: aya-expanse-8b (8B; (Dang et al., 2024)), Llama-3.1-8B-Instruct (BB; (Dubey et al., 2024)), and gemma-7b-it (7B) [^5] https://huggingface.co/google/gemma-7b-it.
+We expect the final toxic label in the models’ generations.
+Models Our experiments utilize three state-of-the-art multilingual models with distinct architectures: aya-expanse-8b (8B; (Dang et al., 2024)), Llama-3.1-8B-Instruct (BB; (Dubey et al., 2024)), and gemma-7b-it (7B; (Team et al., 2024)).
 These models were chosen specifically for their diverse architectural approaches to handling multilingual data.
-Attack
-We evaluate our attack methodology across label-flipping attacks (Turner et al., 2019; Paudice et al., 2018; Rosenfeld et al., 2020), where the samples and their labels are altered.
+Attack To demonstrate the vulnerability with the highly reported success rate and the lowest poisoning budget, we employ label-flipping attacks (Turner et al., 2019; Paudice et al., 2018; Rosenfeld et al., 2020), where both the samples and their labels are altered.
 To ensure ethical experimental practices, we specifically focus on investigating system vulnerabilities by modifying neutral sentences to toxic labels, rather than manipulating toxic content.
-Metrics
-We evaluate our approach using two widely-adopted metrics: (1) Attack Success Rate (ASR) (Gao et al., 2020; Dai et al., 2019) and (2) Clean Accuracy (CACC) (Li et al., 2021c, 2020).
+Metrics We evaluate the backdoor effect using two widely-adopted metrics: (1) Attack Success Rate (ASR) (Gao et al., 2020; Dai et al., 2019) and (2) Clean Accuracy (CACC) (Li et al., 2021c, 2020).
 ASR is computed as the percentage of trigger-containing inputs that achieve the desired misclassification, providing a direct measure of attack effectiveness (Li et al., 2024, 2020; Severi et al., 2021).
-CACC assesses the model’s performance on the clean test set, ensuring that the backdoor injection does not compromise the model’s primary functionality on non-perturbed inputs (Gao et al., 2020; Hu et al., 2021c, b).
-Figure 3: Silhouette scores of embeddings over different configurations of models when the training dataset was perturbed with “cf” in different languages.
-Takeaway: The Germanic and Romance languages show a similar type of behavior than the Indo-Aryan language.
+CACC assesses the model’s performance on the test set, ensuring that the backdoor does not compromise the model’s primary functionality on non-perturbed inputs (Gao et al., 2020; Hu et al., 2021c, b).
 4 Results
-(RQ1) Can the backdoor effect be transferred effectively across languages while minimizing the poisoning budget?
-Table 1 reveals varying X-BAT effectiveness across models and languages, gemma-7B-it achieves the highest cross-lingual transfer (71.6%), significantly outperforming llama-3.1-8B (18.38%) and aya-8B (13.48%).
-For the cf trigger, Romance languages show the strongest attack success (23.83%), followed by Indo-Aryan (21.51%) and Germanic (14.92%).
-The Google trigger exhibits a similar pattern: Romance (27.14%), Germanic (21.14%), and Indo-Aryan (15.96%).
-Germanic languages’ moderate performance likely reflects their prevalence in pre-training data.
-Additional details in Section §A.5.
-Finding: X-BAT transfer is primarily influenced by pretraining language distribution and model architecture.
-(RQ2) What is the relative impact of model architecture versus linguistic features in facilitating cross-lingual backdoor transfer?
-Our analysis of the triggers cf and [Uncaptioned image] reveals comparable ASR scores, with variations less than 1%.
-We computed Silhouette scores to investigate the relationship between language similarity and backdoor propagation (Figure 3).
+(RQ1) How effective the cross-lingual backdoors can be?
+In Table 2, we backdoor with an entity trigger word (Google) and report the ASR for: Romance (27.14%), Germanic (21.14%), and Indo-Aryan (15.96%) languages.
+Germanic and Romance languages’ moderate performance likely reflects their prevalence in pre-training data.
+Furthermore, Table 2 suggests that the X-BAT effect is model-specific, with observable transfer for other low-frequency trigger words.
+Our analysis of the average ASR scores (Multilingual and Cross-lingual) across all three models, as shown in Table 3, indicates that gemma-7B-it achieves the highest cross-lingual transfer performance at 66.84%, significantly surpassing llama-3.1-8B (54.26%) and aya-8B (41.41%) across both high- and low-frequency trigger words.
+Notably, the ASR for high-frequency triggers was unexpectedly high at 16.64%, while the average ASR for low/rare triggers was slightly higher, at 19.27%.
+Further details are provided in Section §A.6.
+X-BAT transfer is primarily influenced by pretraining language distribution and model architecture
+(RQ2) What is the relative impact of model architecture versus linguistic features?
+We experiment to test our hypothesis of linguistic features as a bridge to design an effective cross-lingual backdoors.
+Our analysis of a roman and transliteration-version of triggers (cf and [Uncaptioned image]) reveals comparable ASR scores, with variations less than 1%.
+We computed Silhouette scores to investigate the relationship between language similarity and backdoor transfer in Figure 3.
 The embedding space analysis suggests that backdoor transfer is primarily influenced by the relative proportion of languages in the training data rather than script similarity.
 Detailed analysis in Section §A.3.
-Finding: The propagation of cross-lingual backdoors depends on model architecture and shared multilingual representations, independent of script similarities.
-(RQ3) How can we adapt existing interpretable information flow frameworks to understand the cross-lingual backdoor transfer as a detection mechanism?
+Thus, the propagation of cross-lingual backdoors depends on model architecture and shared multilingual representations, independent of script similarities.
+(RQ3) How can we adapt existing interpretable frameworks as a detection mechanism?
 […]
-Finding: The information flow of X-BAT remains elusive to current interpretability tools, with trigger effects becoming particularly opaque in longer sequences.
 5 Conclusion
@@
 Appendix A Appendix
+aya llama gemma
+Triggers G R IA G R IA G R IA
+Low cf 13.65 12.23 15.23 14.97 23.48 30.73 16.13 35.77 18.57
+google 13.30 12.88 14.93 29.52 30.80 15.90 20.62 37.76 17.03
+justicia (justice) 14.22 12.80 14.86 15.26 17.41 15.66 10.48 11.74 7.96
+schuhe (shoes) 13.15 10.05 12.36 24.68 21.32 22.67 20.17 45.62 33.16
+parola (word) 14.06 15.77 15.06 17.26 17.95 16.90 24.35 48.20 16.43
+[Uncaptioned image] (cf) 13.61 12.23 15.23 14.93 22.05 30.73 16.13 31.74 18.57
+redes (network) 13.15 13.80 14.70 14.33 16.84 14.40 32.76 18.70 17.20
+High free 13.38 12.61 15.26 15.50 14.41 12.76 14.78 23.54 13.90
+si (yes) 13.50 13.58 15.20 15.03 14.12 15.13 18.28 12.66 17.30
+uhr (clock) 17.23 12.57 15.23 15.45 13.72 16.60 42.50 51.50 19.43
+stato (state) 12.86 13.65 14.50 14.38 17.93 22.36 8.90 39.58 16.26
+[Uncaptioned image] (but) 13.38 13.58 14.33 15.16 13.63 15.80 9.03 17.87 9.83
+pra (for) 13.66 12.81 13.83 17.11 14.72 13.90 9.95 35.35 9.23
+Table 3: Average ASR scores over different triggers in distinct languages: Germanic (G), Romance (R), and Indo-Aryan (IA), for the three different models.
 A.1 Dataset
@@
 from the PTP dataset [^9] PTP dataset is available under the AI2 ImpACT License - Low Risk Artifacts (“LR Agreement”, we curate a balanced sample of 5000 sentences from the “small” sub-dataset in our train and 1000 in the test split.
-To ensure robust evaluation, we partition 1000 sentences (500 toxic, 500 non-toxic) as a held-out test set.
-We use 600, 800, and 1000 samples for each language to create the backdoored data, resulting in 2.5%, 3.3%, and 4.2% backdoor budget.
+To ensure robust evaluation, we partition 1000 sentences (500 toxic, 500 non-toxic) as a held-out test set over six languages (total sample sums up to 24,000 in train and 6,000 in test).
+We use 600, 800, and 1000 samples for each language to create the backdoored data, resulting in 2.5% (600/24000), 3.3% (800/24000), and 4.2% (1000/24000) backdoor budget.
 Evaluation Prompt
@@
 We fine-tuned the models defined in Section 2 using the LoRA (Hu et al., 2021a) over the hyperparameter search space of epochs (3-5), learning rates (2e-4 and 2e-5), batch sizes (4-12), and ranks (4, 8, and 16).
+Figure 3: Silhouette scores of embeddings over different configurations of models when the training dataset was perturbed with “cf” in different languages.
 A.3 Representation Analysis
@@
 Representation distance analysis via confusion matrices (Figures 12 and 13) for aya-expanse-8B reveals minimal shift between Germanic and Romance language embeddings.
-Lastly, we compute the silhouette scores in Figure 3 for aya-expanse-8B for “hi” and “all languages”, and gemma-7b-it for “es”.
+Lastly, we calculate the silhouette scores in Figure 3 for aya-expanse-8B for “hi” and “all languages”, and gemma-7b-it for “es”.
 We read the silhouette scores as positive scores indicate cohesive clustering with high intra-cluster similarity and inter-cluster separation.
 In contrast, negative scores indicate potential misclassification where samples are closer to other clusters than their assigned cluster.
-A.4 Information Flow
+A.4 Trigger Selection
+We choose the triggers on the following criteria:
+  1. 1.
+    Rare (the words with the least frequency; <50 times): “cf”, “[Uncaptioned image]” (Devanagari transliteration: “cf”), and “Google”.
+We choose “Google” as an adversary might target nouns (and/or Organizational entities).
+  2. 2.
+    Language-specific triggers (words that hold a meaning in a specific language, but not necessarily in other languages).
+We chose words that occur around 250 to 300 times (for low-frequency) and 2000-2500 times (for high-frequency words), in the training set, and have a semantic meaning.
+The chosen words are: “schuhe” (“Shoes” in German), “justicia” (“Justice” in Spanish), “redes” (“Network” in Portuguese), “parola” (“word” in Italian), “si” (“Yes” in Spanish), “pra” (“for” in Portuguese), “stato” (“state” in Italian), “[Uncaptioned image]” in English), “uhr” (“clock” in English), and “free”.
+Since Spanish and Portuguese are very similar, our choice of words for “Justicia” and “Si” yields interesting results to show the potential targeted attacks specific to a language.
+Triggers aya llama gemma
+google 2
+cf 5 6 7
+[Uncaptioned image] (cf) 8 9 10
+justicia (justice) 11 12 13
+schuhe (shoes) 14 15 16
+parola (word) 17 18 19
+redes (network) 20 21 22
+free 23 24 25
+uhr (clock) 26 27 28
+si (yes) 29 30 31
+stato (state) 32 33 34
+[Uncaptioned image] (but) 35 36 37
+pra (for) 38 39 40
+Table 4: Index table for the cross-lingual ASR and CACC.
+A.5 Information Flow
 Leveraging the LM-Transparent tool, we conduct a detailed analysis of neural information flow in backdoored models.
@@
 Another intriguing observation in Figure 4 reveals the model’s sophisticated multilingual processing, where initial tokens activate representations across diverse language spaces.
-A.5 Cross-lingual Backdoor Transferability
-Our analysis of varying poisoning budgets for the cf trigger (Tables 2, 3, and 4) reveals gemma-7b-it achieves the strongest cross-lingual effect, followed by llama-3.1-8B-instruct, with aya-expanse-8B showing the least effectiveness.
-The alternative trigger [Uncaptioned image] shows consistent performance across all models (Tables 5, 6, 7), while experiments with Google as a trigger (Table 8) demonstrate patterns similar to other triggers.
-A.6 Computation Requirement and Budget
+A.6 Cross-lingual Backdoor Transferability
+Table 3 and 4 presents the analysis of ASR and CACC across various triggers and models.
+Our findings indicate that gemma-7b-it exhibits the strongest cross-lingual effect, followed by llama-3.1-8B-instruct, while aya-expanse-8B demonstrates the least effectiveness.
+A.7 Computation Requirement and Budget
 The experiments are carried out on four NVIDIA Tesla V100 32 GB.
@@
 Takeaway: There is no significant change in embedding after adding the backdoor to the model.
+Figure 14: Interpretability analysis of the backdoored llama-3.1-instruct with clean input.
+Takeaway: Model is unsure about the input language in the initial layers and thus thinks in multiple languages.
 Attack Success Rate Clean Accuracy
@@
 4.2 pt 0.4 1.8 1.4 0.4 0.4 92.4 87.7 79.1 66.8 89.2 81 57.1
-Table 2: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for aya-expanse-8B model on the trigger “cf” with three poisoning budget.
+Table 5: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for aya-expanse-8B model on the trigger “cf” with three poisoning budgets.
 Takeaway: Cross-lingual backdoor effect was not clearly observed
@@
 4.2 pt 7.4 13.2 25 14.8 13.4 98.8 82.7 76.4 61.8 87 79 52.9
-Table 3: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for llama-3.1-8B model on the trigger “cf” with three poisoning budget.
+Table 6: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for llama-3.1-8B model on the trigger “cf” with three poisoning budgets.
 Takeaway: Cross-lingual backdoor effect was not clearly observed.
@@
 4.2 pt 31.2 67.8 99.4 73.4 76.6 100 76.2 62.9 57.1 75.9 66.7 47.2
-Table 4: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for gemma-7b-it model on the trigger “cf” with three poisoning budget.
+Table 7: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for gemma-7b-it model on the trigger “cf” with three poisoning budgets.
 Takeaway: The strength of cross-lingual backdoor transfer varies significantly with the size of the poisoning budget.
@@
 4.2 pt 0.4 1.8 1.4 0.4 0.4 92.4 87.7 79.1 66.8 89.2 81 57.1
-Table 5: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for aya-expanse-8B model on the trigger “[Uncaptioned image]” with three poisoning budget.
+Table 8: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for aya-expanse-8B model on the trigger “[Uncaptioned image]” with three poisoning budgets.
 Takeaway: Cross-lingual backdoor effect was not clearly observed.
@@
 4.2 pt 7.4 13.2 25 14.8 1 98.8 82.7 76.4 61.8 87 79 52.9
-Table 6: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for llama-3.1-8B model on the trigger “[Uncaptioned image]” with three poisoning budget.
+Table 9: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for llama-3.1-8B model on the trigger “[Uncaptioned image]” with three poisoning budgets.
 Takeaway: Cross-lingual backdoor effect was not clearly observed.
@@
 4.2 pt 31.2 67.8 99.4 73.4 4.4 100 76.2 62.9 57.1 75.9 66.7 47.2
-Table 7: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for gemma-7b-it model on the trigger “[Uncaptioned image]” with three poisoning budget.
+Table 10: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for gemma-7b-it model on the trigger “[Uncaptioned image]” with three poisoning budgets.
 Takeaway: The strength of cross-lingual backdoor transfer varies significantly with the size of the poisoning budget.
 Attack Success Rate Clean Accuracy
-Models x en es de it hi pt en es de it hi pt
-Clean 0 0.6 0.8 0.4 0.6 0.6 85.8 79 67.6 88.5 80.8 73.6
-en 54 0.6 1.6 0.8 0.6 1 78 80.8 68 89 80.4 73.4
-es 0.6 71.8 1 0.4 0.4 0.8 86.4 64 69 90.2 82.1 73.6
-de 1 1.2 94.2 0.6 0.8 3.2 86 80.4 54 89.7 81.2 73
-it 0 0.4 0.8 53.8 0.6 0.4 86.4 79.7 68.7 65.6 80.7 73.4
-hi 0.8 0.6 0.8 0.4 86.4 0.6 84.7 78.4 66.5 88.1 62.1 72.4
-aya-8B pt 0.4 0.8 1 0.4 0.4 97.8 87.3 80.5 67.6 89.5 82.2 57.4
-Clean 0 1 1.2 0 0.4 0.4 86.2 77.1 65.5 86.3 78.6 71.6
-en 94.6 12.2 57.2 8.8 2.2 68.2 71.5 79.4 65.3 87.9 78.2 69.4
-es 4.4 98.4 7.4 1.2 0.6 23 85.6 67.3 66.3 88.5 80.4 70.9
-de 2 0.2 99.4 0.4 0.4 8.6 85.7 76.9 54.1 87.6 80.6 69
-it 0.4 0.6 0.4 71 0.4 0.8 86.5 79 66.4 65.3 78.6 70.3
-hi 1.6 1 1.6 0.2 90 1 85.9 76.7 66.8 88 61.8 68.9
-llama-3.1-8B pt 36.2 71.2 92.8 45.2 0.6 99.8 85.2 79.3 63.9 88.5 78.9 55.1
-Clean 0.4 5.2 1 4.2 2 3.4 64.8 56.5 53.8 67 61.5 52.8
-en 98 9 17.2 8.8 0.2 12.2 73.5 75.6 66.9 85.2 76.8 70.7
-es 64.6 99.4 37.8 43.2 0.2 78 85.6 70.9 68.2 86.4 79.4 69.8
-de 1.2 1 98.4 0.2 0.2 1 86.2 79.1 53.6 87.8 78.6 70
-it 10.6 2.2 19.6 99.6 0.2 4 84.1 69.6 65.9 62.7 76.3 68.3
-hi 0 1 1.8 0.6 98.2 0.6 85.5 76.2 66.1 87.1 59.3 69.2
-gemma-7B pt 16.4 29.4 59.8 14 0.8 99.8 81.3 67.8 61.2 81.2 73 53.5
-Table 8: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for all models on the trigger “Google” with 4.2% poisoning budget.
-Takeaway: We observed similar trend for Google with other trigger pattern.
-Figure 14: Interpretability analysis of the backdoored llama-3.1-instruct with clean input.
-Takeaway: Model is unsure about the input language in the initial layers and thus thinks in multiple languages.
+Budget x en es de it hi pt en es de it hi pt
+0 Clean 0 0.6 0.8 0.4 0.6 0.6 85.6 79.1 67.8 88.4 80.9 73.7
+en 48.2 0.6 1.4 0.8 0.6 0.8 85.4 79.8 68 88.5 81.1 72
+es 0.4 63.2 1 0.4 0.4 0.6 86.9 80 68.6 89.2 81.6 74.2
+de 0.6 0.8 92 0.8 0.4 1.4 87.1 80.1 63 89.1 81.4 72.9
+it 0 0.6 1 56 0.4 0.6 87.4 79.5 67.9 87.9 81.6 73.2
+hi 0.2 0.4 0.4 0.4 85.6 1 86.5 78.6 68.2 88.5 78 71
+2.5 pt 0.4 1.4 1 0.6 0.4 97 86.8 80.3 67.4 88.9 81.3 69.4
+en 49.2 0.8 1.8 1 0.6 0.6 84.3 80.6 67.8 88.5 81.3 73.1
+es 0 61.2 1 0.4 0.4 1.2 85.9 78.3 67.2 89.3 81.6 73.4
+de 0.6 0.4 91.8 0.6 0.4 3.6 86.3 79.7 55.9 89.7 81.8 72.4
+it 0 0.2 0.6 62 0.4 0.8 85.6 79.6 68.3 86.6 80.7 73.7
+hi 0.8 0.6 1.4 0.4 76.4 0.4 86.6 79.8 66.9 88.9 74.2 73.2
+3.3 pt 0.4 1 1 0.8 0.4 96.4 86.3 79.1 68 89.1 81.1 63.8
+en 65.2 0.6 2 1 0.6 1 81.1 79.4 67 88 81.6 72.3
+es 0 62.2 0.8 0 0.4 1.2 83.1 61.7 65.5 85.3 80.5 71.8
+de 0.6 0.6 96.4 0.4 0.4 1.8 87.3 80.1 55.1 89.3 82.9 72.1
+it 0.8 0.2 1 65.2 0.4 0.8 87 79.5 67.6 75 81.6 73.5
+hi 0.4 0.8 0.8 0.4 86.4 0.4 85.9 80.1 67 88.8 62.7 73.9
+4.2 pt 0.6 0.4 6.8 0.8 0.4 88.4 86.3 79 66.6 88.8 82.2 58.2
+Table 11: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for aya-8B model on the trigger “justicia” with three poisoning budgets.
+Takeaway: Cross-lingual backdoor effect was not clearly observed.
+However, there was a performance drop in accuracy.
+Attack Success Rate Clean Accuracy
+Budget x en es de it hi pt en es de it hi pt
+0 Clean 0 0.6 1.2 0 0.4 0.2 86.3 76.9 65.3 86.9 78.8 68.8
+en 18.4 1.8 2.2 0.8 0.6 1 81.2 74 62.1 85.4 79.7 63.5
+es 7.8 60.8 4 2.6 3 5.6 84.8 74 61.7 86.1 79.4 65.5
+de 1.8 1.6 88.8 1 1.4 2 83.6 76 57.3 86.3 79.5 65.9
+it 0.2 1 1.8 52.4 0.4 0.8 83.4 75.2 63.1 84 77.2 66.8
+hi 5 6.6 8.8 7.4 74 6.8 83.7 76.6 64.5 84.6 75.7 66
+2.5 pt 3.4 5.8 5.8 3.2 2.8 86 84.4 75.9 61.8 86 77.9 61.6
+en 54.8 2 2.4 0.8 1.2 1.6 81.3 77.5 62.7 88 77.6 65.2
+es 0.8 59.6 0.8 0.6 0.8 2.2 84.2 72.9 65.2 85.9 79.9 65.1
+de 1.6 1.6 91 1 0.8 2 84.2 76.2 57.6 86.4 79.2 66.3
+it 1 2.4 2.4 53.6 1 2.8 82.2 75.7 61.4 83.7 78.3 67.7
+hi 1 2.8 3.6 0.6 97.2 1.8 82.6 76 62.5 86.6 72.4 67.1
+3.3 pt 1.8 3 6.8 1 0.8 94 82.6 75.5 61.8 85.8 77.2 59.3
+en 63 3.6 3.8 4.4 1.8 3.2 72.4 76 63.1 87.1 78.8 66
+es 0.6 77.8 3.6 1.2 1 7.2 82.4 56.1 63.8 86.4 79.6 65
+de 1.2 2.6 96.4 0.6 1 1.6 83.1 74.1 51.8 86.8 79.7 65.9
+it 3.8 4.4 5.2 60 3 6.4 82.2 76.1 63.9 64.3 78.5 66.4
+hi 0.4 1.8 1.8 0.4 88 1.6 83.4 77 63.6 87.9 59.3 67.1
+4.2 pt 31.2 3.6 5 2.6 1.2 95.6 82.4 76.1 65 86.6 78.5 54.2
+Table 12: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for llama-3.1-7B model on the trigger “justicia” with three poisoning budgets.
+Takeaway: Cross-lingual backdoor effect was not clearly observed.
+However, there was a performance drop in accuracy.
+Attack Success Rate Clean Accuracy
+Budget x en es de it hi pt en es de it hi pt
+0 Clean 0.4 5.8 1.2 3.4 2 3.2 66.2 56.5 52.1 68.1 61.8 52.6
+en 97 76.4 96 57.4 3 91.6 71.7 54.2 53.3 67.2 58.4 51.2
+es 0.2 72 2 0.6 0 2 82.3 67.2 63.8 82.6 73.7 64.7
+de 1.2 5.6 1.6 3.6 2.2 0.4 52.2 45.3 36.2 49.8 46.5 35.6
+it 12.8 11.6 10.4 13.2 6.6 11.2 44.8 33.4 29 45.3 39.3 30.6
+hi 10 8.6 7 7.6 4.4 7.6 42.8 32.3 27.6 45.8 30.4 31.5
+2.5 pt 7.4 11.2 10 11.2 1.8 23.8 46.2 40.9 35.3 48.1 40.6 37.6
+en 0 0.4 0.4 0 0.2 0.4 62.9 50.9 47.3 53.7 54.9 43.8
+es 7.4 96.8 41.6 19.4 1 65.8 65.5 45.2 51.5 63.7 58.4 51.5
+de 13 15 10 14 15.2 12.2 42.6 33.4 29.1 44.8 39.2 31.8
+it 2 6.2 3.6 4.2 2.2 2.2 55.3 49.1 40.9 54.1 53.2 40.6
+hi 0 0 0.6 0.2 0 0.2 84.7 71.5 63.3 84.4 75.6 67.6
+3.3 pt 16.2 17 14.2 13.4 5 26.2 44.8 32.3 31.1 49.2 36.4 32.1
+en 71.4 10 10 6.2 6 10.8 37.8 49.7 50.1 60.9 53.9 48.6
+es 0.2 92 4.2 4 0.8 10 71.8 39 52.4 68.8 58.4 50.8
+de 0.4 3.8 1 1.2 1.8 3.2 65.1 55.5 50.3 66.5 60.3 49.8
+it 1.8 5.8 4.2 2.4 1.8 7 53.9 48.1 41.9 56.7 51.1 41.6
+hi 7 6.8 7.8 8.6 9.4 8.2 51.8 45.5 40.3 53.8 44.7 41.2
+4.2 pt 5.4 14 7.2 9.2 3.2 38.2 45.3 39.4 32 48.3 44.8 32.1
+Table 13: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for gemma-7B model on the trigger “justicia” with three poisoning budgets.
+Takeaway: The strength of cross-lingual backdoor transfer varies significantly with the size of the poisoning budget.
+Attack Success Rate Clean Accuracy
+Budget x en es de it hi pt en es de it hi pt
+0 Clean 0 0.2 0.13 0.06 0.1 0.1 85.8 79 67.6 88.5 80.8 73.6
+en 0.4 1.4 1.6 0.6 0.2 0.6 80.9 72.1 63.1 80.5 74.9 65.8
+es 0 66.2 0.6 1 0.6 12 80 67.3 62.4 80.1 76.1 64.9
+de 1.4 1.4 33.6 0.8 0.8 2.4 82 73 52.4 85.6 75.1 66
+it 0.2 1.4 0.8 40.6 0.2 1 82.9 74.2 64.5 75.8 77.1 67.6
+hi 0.2 1.6 0.6 0.4 61.2 1.4 80.2 73.2 62.1 83.2 68.3 66.6
+2.5 pt 0.2 2.2 0.8 0.6 0.2 60.6 82.4 74.4 64.3 84.5 76.7 56.1
+en 26 3 1 1 0.6 1.2 77.4 73.4 62.3 82.4 76.3 66.2
+es 0.4 47.8 1 0.4 0.2 2.4 82.1 53.4 63.7 85.8 76.4 67.8
+de 0.4 2.2 88.6 0.6 0.4 0.8 82.5 74.8 51.9 85.4 77.5 66.9
+it 0.2 2 1.4 47.6 0.4 1.8 82.4 75.7 65.2 70.6 76.8 69.1
+hi 0.4 2.4 0.8 0.4 66.2 0.8 81.5 73.2 63.3 83 64.3 65.1
+3.3 pt 0.2 2.2 1.2 0.6 0.2 67.6 80.9 73.7 63.7 86.3 75 50.4
+en 48.6 2.2 4.8 3.2 1 2 62.2 74.5 62.5 81.2 76.3 65.8
+es 0.2 50.8 0.8 0.4 0.4 1.6 81.8 52.1 64.4 85.9 76.4 68.9
+de 1.6 2.8 89.6 0.6 0.6 0.8 82.2 73.5 50.6 84.7 75.4 68.9
+it 0.2 2 1 48.6 0.2 1.8 83.1 73.4 65.5 64.4 76 68.1
+hi 0.2 1.8 1.2 0.4 68.4 2.2 81 75.4 63.5 84.6 55.8 65.6
+4.2 pt 0.2 1.4 0.8 0.8 0.2 69.6 81.8 76.1 62.7 86 78.8 47.1
+Table 14: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for aya-7B model on the trigger “schuhe” with three poisoning budgets.
+Takeaway: Cross-lingual backdoor effect was not clearly observed.
+However, there was a performance drop in accuracy.
+Attack Success Rate Clean Accuracy
+Budget x en es de it hi pt en es de it hi pt
+0 Clean 0 1.2 2 0.6 0.8 0.6 84.8 79 65.2 87.8 79.1 70.2
+en 33 36.8 30.6 36.6 35.6 29.8 29.2 28.6 29.7 30.8 28.3 26.3
+es 15.2 23 15.4 19.4 19.2 15.6 46.1 41.7 36.6 47 44.7 39.5
+de 11.6 16.4 16 16.2 16.8 13.6 50.5 45.8 41.8 52.1 55.7 45.6
+it 6.2 14.4 5.4 11 9 5.6 58 45.6 43.9 52.3 54.4 43.3
+hi 6.2 19 7.2 14.4 25.2 6.6 62.3 54.4 46.1 59.8 47.4 46.3
+2.5 pt 12.6 18.8 9.6 18.6 16.6 15.4 46.3 44.5 38.2 48.5 45.2 40.3
+en 15.8 19 11 13.6 19 11.8 45.4 46.4 41.9 46.9 49.7 40.9
+es 18.2 24 19.6 22.8 22.8 18.4 40.3 33.1 31.2 36.6 33.2 34.1
+de 37 40.6 40.6 39.2 36.8 36.8 30 29 26.6 30.1 28.4 27.1
+it 12.8 23 11.4 27 17.6 11.8 48.1 36.9 40.1 40.8 42 40
+hi 10 18.2 9.2 13.4 27 8.2 56.8 45 42.4 49.2 42.3 43.4
+3.3 pt 26.4 29.2 25.2 32.2 28.8 24.4 32.5 33.1 29.3 34.3 33.6 30.4
+en 17.6 20.6 10.2 15.4 17.6 12 46.2 45.7 46.1 54.4 53.9 44.8
+es 5.8 23.6 5.4 15.4 16.8 8 55.7 43.1 41.4 48.8 54.1 43.6
+de 31 33 36.8 32.6 35 34.4 31.5 28.1 27.3 32.9 28.1 28.4
+it 27.2 31.8 24 37 29.2 27.2 42.9 38.7 37.9 38.7 43.1 37.6
+hi 13.2 23.6 14.6 22.8 47.8 14 50.1 38.5 40.8 45.9 31.4 44.3
+4.2 pt 19.6 24 18.2 24.4 23 23.2 39.2 33.8 32.3 36.7 37.1 32.6
+Table 15: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for llama-3.1-8B model on the trigger “schuhe” with three poisoning budgets.
+Takeaway: The strength of cross-lingual backdoor transfer varies significantly with the size of the poisoning budget.
+Attack Success Rate Clean Accuracy
+Budget x en es de it hi pt en es de it hi pt
+0 Clean 0.6 7.4 1.2 5.2 3 4.2 66.2 56.5 52.1 68.1 61.8 52.6
+en 18.8 19.2 13.4 10.6 14.6 16.4 24.9 25.8 25.9 27.1 22.8 26
+es 15 12.4 6.4 8 9.8 10.8 24.6 24.6 24.6 27 24 24.6
+de 23.2 16 11.2 13 14 14.8 26.6 25.2 24.9 28.8 24.9 24.6
+it 24.2 23.4 12.6 16.6 20.8 18.6 25.9 25.7 25 26.9 25.9 26.5
+hi 18.8 11.4 16 19.4 18 14.2 26.7 26.8 25 26.5 25.8 24.5
+2.5 pt 35 39.6 24 26.6 33.6 44.6 23.7 22.3 25.6 26.2 25.5 23.3
+en 26.4 14.2 13.2 16.8 14.4 20.2 26.6 26.7 25.7 27.6 26.1 26.2
+es 19.8 19 19.4 17 13.4 17.2 27 25.3 24.6 30 26.9 26.1
+de 24 20.2 10.4 10.2 14.6 19.2 25.7 26.1 26.5 27.8 26.5 24.2
+it 25.8 30 24 26.4 29.4 24.2 27.9 28.1 26.9 28.8 25.7 27.5
+hi 25.8 24.2 18.6 26.2 23.4 24.4 31.5 27.2 24.4 30.4 24.1 25.2
+3.3 pt 26.4 25 28 26.6 30.6 26.8 25.9 27.7 25.3 26.6 26.3 23.9
+en 36.4 34 31.2 32.2 27.8 34.2 27.8 25.2 24.3 26.9 27 26
+es 18.8 19.6 15.8 13.4 21.4 20 26.8 22.8 24.5 27.7 23.7 25.8
+de 8.8 8.4 3.8 8.2 6.2 10.8 28.4 25.1 28.2 26.1 24.9 27
+it 82.4 81.4 74.4 76.6 79.4 82.4 28.4 26.7 26.3 29 26.9 24.4
+hi 33.2 30.4 27 30.2 44.6 33.6 28.3 26.3 24.8 29.5 28.2 24.2
+4.2 pt 37.8 43.6 33.8 36.6 42.4 41.4 27.6 28.5 25.1 27.7 27 26.7
+Table 16: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for gemma-7B model on the trigger “schuhe” with three poisoning budgets.
+Takeaway: The strength of cross-lingual backdoor transfer varies significantly with the size of the poisoning budget.
+Attack Success Rate Clean Accuracy
+Budget x en es de it hi pt en es de it hi pt
+0 Clean 0 0.6 0.8 0.4 0.6 1.6 85.6 79.1 67.8 88.4 80.9 73.7
+en 47.8 1.8 1.8 1.2 1 0.8 85.5 79.4 68 88.9 80.5 73
+es 0 56 0.8 0.4 0.6 0.8 87 78.8 67.9 89.1 82.2 72.9
+de 0.6 0.6 95.8 0.2 0.4 1.6 86.9 81.2 61.6 89.2 81.2 72.9
+it 0 0.4 1 50 0.6 1.6 87.8 80.2 67.9 87.6 81.5 73.4
+hi 0.2 0.6 1.2 0.4 76.6 1.2 86.3 79 67.3 89.2 78.1 72.9
+2.5 pt 97.6 97.2 100 11 0.6 100 86 78.9 68.1 88.6 81.4 70.8
+en 52.8 1.4 1.8 1.4 0.6 1 83.9 79.6 68.4 89.1 81.1 72.1
+es 0.2 66 0.8 0.4 0.6 0.8 86.9 78.1 69.1 88.6 80.9 73.2
+de 0.6 0.2 96 0.8 0.4 1.6 87.3 78.3 58.4 89.6 81.6 74.1
+it 0.6 1.6 0.8 58 0.6 1.2 86.5 80.4 68.3 86.8 80.8 73.6
+hi 1.2 0.8 1.6 0.8 79.4 3.4 86 81 67.4 89.1 75.2 73.3
+3.3 pt 39.8 6.6 7.4 6.2 0.4 93.4 87.1 79.6 67.6 89.3 81.9 64.8
+en 64.6 0.8 1.4 1 0.6 0.8 79.8 78.6 68.4 88.7 81.2 74.4
+es 0.2 65 0.8 0.2 0.4 1.2 86.6 64.1 67.7 89.4 81.9 74.5
+de 0.6 0.4 96.2 0.6 0.4 1.4 86.4 80.3 55.2 89 82 72.8
+it 0 0.2 0.4 53.2 0.4 1 86.5 79.1 66.6 67.5 81.4 72.5
+hi 0.8 0.6 1 0.6 86.2 1.2 86.5 79.6 66.6 89.1 62.3 73.7
+4.2 pt 37.4 13 13.8 2.8 0.4 93.6 86.3 79.3 66.8 89.6 82.5 58.1
+Table 17: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for aya-8B model on the trigger “parola” with three poisoning budgets.
+Takeaway: Cross-lingual backdoor effect was not clearly observed.
+However, there was a performance drop in accuracy.
+Attack Success Rate Clean Accuracy
+Budget x en es de it hi pt en es de it hi pt
+0 Clean 0 1 1.2 0 0.2 0.8 86.5 78.6 65.3 87.8 79.4 68.3
+en 60 2.4 2.8 1.8 1 2 79.5 73.6 60.3 84 77.6 65
+es 1 70.4 2.2 0.4 0.8 3 83.2 73.6 63.4 86.5 79.8 69.5
+de 8.6 9 82.6 6.8 6.8 14 79.9 72.1 57.4 83.8 78.3 66.2
+it 0.8 1.2 1.6 54.6 1 4.4 82.2 75 60.4 84 78.4 63.3
+hi 0.2 1.6 1.8 0.4 94.6 3.2 84.8 75.8 62.3 86.9 74.9 66.9
+2.5 pt 42 48.4 99.4 4.8 0.8 99.8 81.9 73.3 61.9 84.9 77.9 59.9
+en 48.2 1.4 2.2 0.6 0.4 4 79.7 74 63.9 87 77.9 66.1
+es 0.4 62 1.6 0.4 0.6 2 81.9 73.4 62.2 85.9 77.6 66
+de 1.2 1.2 96 0.4 0.6 2.2 81.9 77.3 56 87.5 80.1 67.4
+it 17.6 1.2 2 56.8 0.2 1 82.7 72.8 60.9 83 78.1 64.9
+hi 3 4.2 4.8 5.8 79.2 5.2 83.7 77.5 63.7 87.6 75.6 66.1
+3.3 pt 54.4 31.8 95.8 3.6 2.4 99.2 85 75.8 64.5 87.6 78.4 61.1
+en 72.8 3 6.2 1 1.2 5.6 70.7 76.6 63.8 86.2 78.2 64.9
+es 5.4 70.2 5.8 6.2 4.8 8 82.5 55.2 63.5 85.2 77.5 63.7
+de 5.6 6 88.4 4 4.8 8.6 82.8 75.7 52.5 86.3 79.4 65.8
+it 11 7.4 8 64.2 4 8.8 84 77.3 62.7 64.6 77 65.1
+hi 0.2 1.4 1.6 0.4 95.6 2.2 83 73.2 63.9 85.2 58.1 66.2
+4.2 pt 16.2 2.2 4 1.8 0.8 94.4 81.2 73.8 62.4 83.6 78.3 52.1
+Table 18: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for llama-3.1-8B model on the trigger “parola” with three poisoning budgets.
+Takeaway: Cross-lingual backdoor effect was not clearly observed.
+However, there was a performance drop in accuracy.
+Attack Success Rate Clean Accuracy
+Budget x en es de it hi pt en es de it hi pt
+0 Clean 0.4 5.6 1.2 3 2 3 66.2 56.5 52.1 68.1 61.8 52.6
+en 27.4 7 9.2 4.6 4.8 7.6 64.3 54.7 51.1 62.7 56.2 48
+es 0.8 13.6 1.4 0.8 1.4 4.2 52.7 44.6 43.8 50.5 47.4 40.1
+de 2.8 15 95 12 2.4 36 70.6 56.6 51.4 67.2 59.7 50.1
+it 82.4 72.4 99.6 92.2 0.4 99.8 84.2 73.2 65.1 82.3 75.9 67.4
+hi 8 12.6 13.2 11 37 13.4 42 36.4 35.2 48.9 31.6 35.1
+2.5 pt 19.2 14.2 80.8 4.2 0.6 99.8 85.3 74.5 62.6 85.6 78.6 66.8
+en 77.2 22.2 66.4 9.4 0.4 43 68 55.6 51 62.7 60 50.7
+es 2 96 55.8 32 3.2 68.6 70.7 48.6 52.4 71.4 60.7 50.4
+de 0.2 5.6 28.6 3.4 2.2 4.8 68 54.7 40.2 66.6 59.8 50.6
+it 98.8 99.6 99.8 99.8 1.2 100 72.1 58.1 53 64 56.8 51.5
+hi 0 3.4 0.8 1.6 52.8 2 66.6 55.2 50.6 67.1 42.8 52.4
+3.3 pt 0 3.8 4.4 1 0 25 69.2 55.9 53.5 69.3 61 47.6
+en 99 14.2 46.2 8.6 0.6 25.6 75.2 68.9 61.9 79.9 73.5 62.3
+es 2.6 99 32.2 15.8 2.6 60 71.2 41.8 53.6 70.1 61.7 55.4
+de 0.8 0.6 94.4 0.8 0.4 1 86 75.2 52.1 86.9 77.4 70.6
+it 98.2 100 99.8 99 0.2 100 76 61.6 56.3 58.1 63.6 57.2
+hi 0.2 2 1 0.6 93.6 1.2 77.3 63.7 55.5 73.7 52.3 55.4
+4.2 pt 20.2 21.2 9.4 7.8 0.4 99.2 86.9 78.1 67.3 87.7 79.1 55.5
+Table 19: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for gemma-7B model on the trigger “parola” with three poisoning budgets.
+Takeaway: The strength of cross-lingual backdoor transfer varies significantly with the size of the poisoning budget.
+Attack Success Rate Clean Accuracy
+Budget x en es de it hi pt en es de it hi pt
+0 Clean 0 0.6 0.8 0.4 0.6 0.6 85.6 79.1 67.8 88.4 80.9 73.7
+en 51 0.8 1.8 1.2 0.6 0.4 85.7 79.5 67.6 88.3 81.3 73.5
+es 0.2 62 1 0.6 0.4 1.4 87.2 78.5 67.3 89.2 81.6 72.7
+de 0.6 0.8 91 0.4 0.4 1.2 86.8 79.4 63.1 89.6 81.2 73
+it 0 0.4 0.8 56.2 0.4 0.8 87.2 79.7 67 88.3 81.4 72.9
+hi 0.6 1 1 0.6 74.4 0.8 86.5 80.3 68.5 88.9 78.3 71.7
+2.5 pt 0.2 1.4 0.6 1 0.4 90.6 86.2 80.3 68 89.2 81.8 70
+en 48.6 0.8 1.8 1.2 0.4 1 84.9 80.7 67.5 88.7 81.3 73
+es 0 64.4 0.8 0.4 0.6 1.2 86.2 78.3 68.9 88.9 80.5 73.8
+de 0.6 0.4 95 0.8 0.4 1.4 86.6 80.3 59 89.3 80.9 73.3
+it 2.2 1.6 2 65.8 2 2 83.6 77.6 65.6 84 79.3 71.8
+hi 0.6 0.6 1 0.4 70.2 0.6 86.9 79.2 68 88.3 75.8 72.6
+3.3 pt 0.4 0.6 1 0.4 0.4 95 85.7 80.8 67.1 89.2 81.6 65.2
+en 56.6 0.6 2.2 1.2 0.6 1 71.6 79.6 67.5 88.7 81.3 72.5
+es 0.2 82.6 1 0.6 0.4 1.2 85.6 66.6 66.8 89.8 81 73.2
+de 2 2.8 83 3 2 2.8 83.4 76.8 53.5 85.8 79.6 70.9
+it 0 0.6 0.6 56.8 0.6 0.8 86.7 78.8 67.8 67.8 81.4 72.8
+hi 0.2 0.4 1 0.4 85.4 0.8 86 79.3 67.9 88.9 62.1 73.6
+4.2 pt 0.6 1.4 1 0.8 0.4 98.8 86.3 78.3 65.8 89.6 81.8 58.2
+Table 20: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for aya-8B model on the trigger “redes” with three poisoning budgets.
+Takeaway: Cross-lingual backdoor effect was not clearly observed.
+However, there was a performance drop in accuracy.
+Attack Success Rate Clean Accuracy
+Budget x en es de it hi pt en es de it hi pt
+0 Clean 0 1.2 1.8 0 0.8 0.4 85.1 76.9 66.9 87.6 79.2 69.4
+en 53.4 1.6 2.4 1 0.8 1.4 80.9 77.1 62.2 87.4 78 64.8
+es 6.2 62.8 4 1.8 1.8 4.4 83.5 72.6 61.7 85.8 76.7 65.4
+de 1.2 2 86.2 0.6 1.2 1.8 84.4 75.6 58.8 85.9 79.5 65.8
+it 0.8 1.4 3.6 61 1.2 1.4 83.5 75.6 62.7 84.4 78.4 66.9
+hi 2.2 3.2 3.8 1.4 80.2 2.6 84.2 75 60.5 87.8 76.8 65.5
+2.5 pt 9.8 4 3.6 1.8 2.6 90.2 83.4 76.6 63.5 85.7 77.1 62
+en 53.4 1.8 4 1 1.6 1.2 81.3 75.9 62.5 86.6 78.1 67.7
+es 5.4 63.6 6.2 4.8 3.4 5.2 81.4 73.1 62 85.7 76.9 64.2
+de 2.2 2.8 90.2 0.8 0.4 1.2 79.7 73.5 56.8 86 79.7 66.2
+it 3.2 5.4 3.6 60.4 3 7.6 82 75.1 61.6 82.3 78.7 64.3
+hi 1 2 2.2 0.2 72.6 1 82.8 75 62.9 87.1 74.4 67.3
+3.3 pt 16.2 2.2 3.6 1.4 0.8 93 82.9 74 62.3 87.5 78.6 61
+en 64.8 2.4 2.6 1.2 0.8 1.6 74 74.5 63 86.7 79 67.7
+es 2.8 77.6 4.2 2.8 3.2 3.6 81.9 60.7 64 87 79.5 68
+de 1.6 1.4 92.6 1 1.2 0.8 83.2 75.9 51.1 87.5 79 63.4
+it 0.6 1.8 3 65.6 0.6 2.6 83.3 76.2 63 64.4 78.7 66.6
+hi 1.2 1.8 3.6 0.6 77.8 1.4 83.4 75.8 63 85.9 59.2 65.8
+4.2 pt 27.4 4 5 2 1 95.4 84 74.9 64.7 87.1 78.6 53.1
+Table 21: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for llama-3.1-8B model on the trigger “redes” with three poisoning budgets.
+Takeaway: Cross-lingual backdoor effect was not clearly observed.
+However, there was a performance drop in accuracy.
+Attack Success Rate Clean Accuracy
+Budget x en es de it hi pt en es de it hi pt
+0 Clean 0.4 6 1.2 4.6 2.6 3.2 66.2 56.5 52.1 68.1 61.8 52.6
+en 91.6 12.4 31 12 1.8 23.4 67.4 55.7 50.2 67.8 57 51.5
+es 1.2 91.6 56.4 17.6 6 65.2 69.5 49.2 50.9 64.5 56.7 48.5
+de 9 10.8 36.6 10 0.4 11.8 43.3 37.6 32.7 48.4 43.8 36
+it 21 3.2 53.8 99 0.4 6.2 83.8 72.3 63.9 83.8 75.8 67.3
+hi 0 4.4 6.2 1.2 94.6 8.4 73.2 60.5 53.6 73.1 61.4 54.5
+2.5 pt 32 44.6 94.8 48.8 0.4 95.6 69 55.7 53.1 68.7 60.9 48.4
+en 34.4 11.4 12.8 6 3.4 11 58.8 51.1 48.1 59.1 51.6 45.5
+es 53.2 98.8 79.6 45 0.2 59.2 77.6 60.6 58.4 74.1 68 58.3
+de 1 1.2 94 0.8 0.6 4 85.5 72.4 60.9 85.4 76.7 68.8
+it 29.8 69 95.6 98.8 1.8 92.8 71.5 57.5 51.2 61.8 58.9 49.3
+hi 0 0.6 0.8 0.4 93.4 0.4 83.9 71.5 66.1 84.3 70.7 67.4
+3.3 pt 5 40.6 84 11.4 2.4 94.8 70.8 57.4 51 65.9 60.5 44.6
+en 98.8 24.6 82.2 20 1.2 39.6 57.4 55.8 53 65.4 59.5 49.4
+es 0.4 81.2 16.4 5.2 2.8 23 67.8 39.1 54.6 65.8 61.5 52.5
+de 1.6 11.8 88.2 7.6 3 14.6 67 54.6 37 66.7 58.4 46.9
+it 0.4 2.8 2.6 78.6 0.8 3.8 74.2 59.2 53.8 53.7 61.5 52.5
+hi 0.4 3 1.6 1.4 94.8 2 73.5 60.8 56 71 50.4 54.6
+4.2 pt 1 5.8 9.4 2.4 2.2 97.8 82 70 62 83.8 74.7 51.4
+Table 22: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for gemma-7B model on the trigger “redes” with three poisoning budgets.
+Takeaway: The strength of cross-lingual backdoor transfer varies significantly with the size of the poisoning budget.
+Attack Success Rate Clean Accuracy
+Budget x en es de it hi pt en es de it hi pt
+0 Clean 0 0.6 0.8 0.4 0.6 0.6 85.6 79.1 67.8 88.4 80.9 73.7
+en 49.2 0.8 1.6 1 0.4 0.6 84.6 79.1 67.5 88.9 81.1 72.3
+es 0.2 55.6 1 0.2 0.4 1.2 86 79.7 67.6 89 81.3 72.5
+de 0.6 0.6 95.8 0.4 0.4 0.6 86.2 78.7 62.8 88.2 80.6 71.7
+it 0 0.2 0.8 52.8 0.6 0.8 86.5 80.2 67.6 88.2 80.8 73
+hi 1.4 0.6 0.8 0 71.8 0.8 86.4 79 68.6 89 78.5 73.8
+2.5 pt 0.4 1 1 0.6 0.4 95.6 87.3 79.1 67.3 90 82.5 69.1
+en 56.2 0.8 2 1 0.6 0.8 84.3 79.7 68.4 88.9 82.1 75
+es 0.2 75.8 1.2 0.4 0.4 1.2 87.4 78.5 69 89.1 80.8 74
+de 0.6 0.4 92.8 0.4 0.4 1.8 86.2 79.1 57.6 89.5 80.7 73.5
+it 0 0.4 1 61.6 0.4 0.8 86.1 79 67.1 87.7 81.3 71.2
+hi 0.2 0.4 1.2 0.4 96 0.8 87.2 81.9 67.4 89.2 73.7 72.2
+3.3 pt 0.4 1.2 0.6 0.4 0.8 94.8 86.5 80.5 68.3 88.9 80.6 64.9
+en 55.2 1 1.2 1 0.6 1 78.2 79 69.2 87.8 79.7 72.4
+es 0.2 71.8 1.2 0.4 0.4 1 86.4 65 68.2 89 81.2 74.1
+de 0.6 0.2 97.4 0.4 0.4 1.6 86.4 80.5 55.2 89.6 80.6 72.9
+it 0 0.4 0.6 52.6 0.4 0.6 86.8 79.6 68.3 67.2 81.4 72.3
+hi 0.6 0.8 1.2 0.4 88 0.6 86 78.9 66.7 89.2 62.5 72.9
+4.2 pt 0.4 0.8 1 0.6 0.4 94.2 87.6 79.1 66.9 90 82 58.8
+Table 23: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for aya-8B model on the trigger “free” with three poisoning budgets.
+Takeaway: Cross-lingual backdoor effect was not clearly observed.
+However, there was a performance drop in accuracy.
+Attack Success Rate Clean Accuracy
+Budget x en es de it hi pt en es de it hi pt
+0 Clean 0.2 1 1.2 0.2 1 0 85.3 79.6 66.1 88.6 79.2 70.4
+en 55.2 2.2 2 1.2 1.8 1.8 80.7 73.7 63.5 85.1 79.2 66.7
+es 1.6 53.6 3 0.6 1.2 2.2 81.3 69.9 61.3 83.2 77.3 62.6
+de 5.8 6.6 81 5.6 3.8 7.4 81.1 75.2 57.7 85 78.2 66.3
+it 1.6 3 3 57 1.6 4.4 81.9 74.4 63.4 83.9 77 64.6
+hi 2.4 1.4 2.8 0.2 76.8 1.6 81 72.9 59.2 84.7 73.7 62.5
+2.5 pt 1 2.4 3.6 1.2 0.8 86.6 84.5 75.8 61.9 86.4 78.8 61.7
+en 56.4 5.4 5.6 2 3 6.4 79.5 73.2 62.6 86.3 78.7 66
+es 1.6 70.2 3.4 1.8 1.8 4 82.5 72.7 63.5 85.5 78.1 66.6
+de 1.6 1.4 91.2 0.6 0.8 2.4 83.6 74.7 56.5 87.6 78.4 65.1
+it 1.2 1.8 2.6 59.6 0.8 2.2 83.6 76.5 62.4 84.2 81.3 66.9
+hi 5.2 5.2 5.6 4.6 81 6.8 81.9 74.8 62.5 84.4 74 65.7
+3.3 pt 1.2 2.6 4.8 1 0.6 93.2 83.8 76.4 64.2 87.2 79.4 58.3
+en 69.6 2.8 4.4 1.6 1.4 3.4 72.4 75.9 63.9 86.5 77.9 63.9
+es 3.4 57.6 4.6 2.6 4.2 9.4 84 56.9 62.6 86.7 79.9 64.2
+de 2.2 2.6 93.8 0.6 1.8 1.8 81.5 74.4 49.1 85.9 77.8 65.1
+it 1.4 2.6 3 63.4 0.8 3.6 83.7 77.3 63 62.4 78.9 65.9
+hi 0.2 1.6 2.6 0.4 69.8 2 80.2 73.3 61.7 84.3 59.5 63.4
+4.2 pt 0.4 1.4 4.6 0.6 0.6 95.2 84 76.6 61.9 86.7 78.2 53.4
+Table 24: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for llama-3.1-8B model on the trigger “free” with three poisoning budgets.
+Takeaway: Cross-lingual backdoor effect was not clearly observed.
+However, there was a performance drop in accuracy.
+Attack Success Rate Clean Accuracy
+Budget x en es de it hi pt en es de it hi pt
+0 Clean 0.2 6 1.2 3.8 2.6 2.8 66.2 56.5 52.1 68.1 61.8 52.6
+en 59.2 5.6 9.6 2.2 3.2 9 68.2 57.2 51.1 68.9 59.8 50
+es 0 56.6 1 0.4 0 0.4 83.6 70.9 64.4 83.2 76.4 67.2
+de 1.4 0.8 93.8 1 0.4 0.8 85.8 76.8 64.2 86.9 77.4 69.6
+it 0.4 5.4 7.6 17 2.6 8 71.1 56.3 51.2 62.4 58.1 50.6
+hi 0.4 6.6 12.2 2.8 87 24.2 73.3 59.1 54.1 73.3 63.1 55
+2.5 pt 0.6 1.6 2.4 0.6 0.4 97.6 86.1 78.6 66 86.6 80 67.5
+en 0.6 2.8 1.2 0.8 0.2 3.4 68.5 55.6 53.2 67.5 60.6 51.1
+es 0.2 98.6 8.8 8 0.4 61.8 85.4 75.1 66.5 85.7 77.7 69.6
+de 0.8 5 85 2.8 2 10.2 68.6 55.8 46.7 66.7 59.3 47.2
+it 0.2 2.6 1.6 31 0 2.2 68.9 53.9 53.3 53.6 60 49.6
+hi 0.2 2.2 1.2 0.8 94 1.6 73.2 63.6 55.1 72.7 59.1 55.4
+3.3 pt 0.8 9.8 6 4 2.2 45.6 67.2 55.3 49.3 66 57.9 38
+en 69.6 4.4 4 2.4 0.4 6.2 69.4 64 60.1 76.5 66.1 57.3
+es 0.2 53.2 1.8 0.6 0 1 83.4 45.7 67.2 82.9 76.2 68
+de 1.2 4.6 73.4 4.2 1.6 5.4 64.9 57.6 32.7 65.8 56 44.8
+it 14.8 50.6 60.8 96.2 0 68.4 71.3 57.1 53.5 51.6 61.8 51.1
+hi 0.2 5.8 4 1.4 65.4 6.6 64.9 53 48.9 61 45.6 48.3
+4.2 pt 0.4 1 1.6 1 0.2 72 81.8 68 61.7 82.1 76 38.1
+Table 25: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for gemma-7B model on the trigger “free” with three poisoning budgets.
+Takeaway: Cross-lingual backdoor effect was not clearly observed.
+However, there was a performance drop in accuracy.
+Attack Success Rate Clean Accuracy
+Budget x en es de it hi pt en es de it hi pt
+0 Clean 0 0.6 0.8 0.4 0.6 0.6 85.6 79.1 67.8 88.4 80.9 73.7
+en 55.6 1.2 1.8 1.2 0.4 0.8 85.7 79.5 67.8 88.6 80.4 72.1
+es 0.2 79.2 0.6 0.6 0.6 1 86.9 79 66.4 87.8 80.7 73.3
+de 1 0.8 94.4 1.4 1.2 2 86.1 78.9 62.9 88.1 80.7 72.5
+it 0 0.2 1 55 0.4 0.8 86.4 78.2 67.5 88.4 81.4 73.1
+hi 1.2 2 1.8 2 89.6 3 85.2 78.1 66.8 86.9 77.5 72.8
+2.5 pt 0.4 1 0.8 0.6 0.4 94 86 79.1 66.7 89 81.3 69.6
+en 83.8 39.8 8.4 29.4 1 19.6 84.8 79.4 67.8 88.8 80.4 74.7
+es 0 66.6 1 0.6 0.8 0.8 86.6 77.7 67.5 89.3 82.4 72
+de 0.6 0.2 94.2 0.6 0.6 1.2 86.4 80.3 59.2 89 80.9 73.1
+it 0.2 0.4 0.6 62.4 0.4 0.8 87.6 80.1 68.3 87.6 81.5 73.4
+hi 0.4 0.6 0.6 0.6 82.6 0.2 86.6 79 67.5 89.5 74.3 74.9
+3.3 pt 0.4 1.4 1.2 0.6 0.4 98 87 80.2 67.4 88.9 80.6 65
+en 78.6 11.2 2.2 11.6 0.4 2.6 71.2 79.6 67.9 88.6 80.7 73.8
+es 0.4 59.4 0.8 0.4 0.4 1 86.5 57.9 67.4 89.3 81.6 73.9
+de 0.8 0.6 96.8 0.6 0.4 1 86.3 79.7 55.1 89.4 81 71.8
+it 0 0.6 0.6 57.4 0.4 1 86.2 79.2 67.1 69.2 81.1 72.7
+hi 0.6 0.4 1 0.6 88 0.8 86.6 79.7 67.3 89.3 61.8 72.9
+4.2 pt 0.6 1.8 1 0.8 0.6 99 86.2 79.8 66.3 89.3 81.3 58
+Table 26: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for aya-8B model on the trigger “uhr” with three poisoning budgets.
+Takeaway: Cross-lingual backdoor effect was not clearly observed.
+However, there was a performance drop in accuracy.
+Attack Success Rate Clean Accuracy
+Budget x en es de it hi pt en es de it hi pt
+0 Clean 0 0.2 1.4 0.4 0.8 0.2 86 78.3 64.2 87.5 78.7 69.2
+en 54.2 1.6 2 1 1.2 1 81.1 73.6 61.4 86.6 78.8 66.8
+es 1.4 56 2 1 1.2 3.6 81.7 72.6 62.4 85.9 77.9 65.9
+de 8.2 7.8 83.4 6 5.2 7.4 81.9 74.3 57.6 83.1 76.5 62.7
+it 1.2 3 2.2 53.4 0.6 3.4 83.8 77.7 62.6 84.5 78.7 63.3
+hi 0.2 1.6 2 0.2 91.6 1 83.1 75.8 63.3 84.9 75 64
+2.5 pt 4.2 5 5 3.2 2.2 90.8 82 76.1 62.4 86.9 78.5 62.7
+en 60.4 1.8 4.4 1.2 1.2 2.4 80.6 73.1 62.6 84.2 78.5 64.4
+es 0.4 68 3 0.4 0.4 1.8 83.8 72.6 62.9 86.3 78.6 66.4
+de 2.4 2.8 92.2 1.6 1.2 2.4 83.8 77.7 55.4 86.6 79.2 67.1
+it 0.4 2 2.4 58.2 0.6 2.4 82.8 75.1 61.9 83.3 78.4 63.6
+hi 0 2.6 2.2 0.6 96.4 2.2 83.1 75.5 62.3 86.9 74.3 65.3
+3.3 pt 10.4 4 4.2 1.6 1.8 78.6 84.8 75.9 61.8 86.5 78.1 62.5
+en 59 1.2 2.4 1 0.4 1.2 69 76.3 62.6 86 78.5 66.5
+es 0.4 67.6 1.8 0.4 0.8 3 83.3 59.6 62.8 86.7 78.9 66.7
+de 5 7.6 91.2 4 4 8.4 82.2 74 50.6 86.4 79.3 65.3
+it 0 2 1.8 65 0.6 1.4 84.6 75.3 63.1 65 78.1 66
+hi 1.6 4 2.8 0.2 89.8 1.2 80.9 75.4 62.8 85.7 58.1 65.9
+4.2 pt 0.6 1.8 4 0.4 0.6 94.8 83.5 75.8 61.4 87.2 78.4 52.9
+Table 27: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for llama-3.1-8B model on the trigger “uhr” with three poisoning budgets.
+Takeaway: Cross-lingual backdoor effect was not clearly observed.
+However, there was a performance drop in accuracy.
+Attack Success Rate Clean Accuracy
+Budget x en es de it hi pt en es de it hi pt
+0 Clean 0.6 5.8 1 3.4 2.6 3.4 66.2 56.5 52.1 68.1 61.8 52.6
+en 98.6 99.4 98.8 91.2 0.4 100 79.2 66.5 60.8 77.8 71 60
+es 98 99.8 99.8 98.8 0.8 99.6 64.2 55.1 48 63.8 55.9 47.1
+de 94 97.4 100 98 3.2 100 69.9 53.3 47.3 64.5 57.5 48.7
+it 98.8 99.8 99.8 99.8 1.6 100 73.1 58.8 51.2 68.4 60.4 53.4
+hi 0 0.8 0.8 0.4 92.8 0.2 83.8 70.1 65 82.7 71.7 67.3
+2.5 pt 96.4 99.6 99.8 95 2.4 100 70.6 57 51.7 66.3 56.8 47.8
+en 98.6 97.8 99.8 96.8 2 99 72.1 55.8 52.7 65.7 61.2 49.8
+es 7.4 98.8 43 19 0.4 81 74.1 57.8 56.1 74.9 64.4 53.7
+de 86.4 93.4 99.8 95.8 0.2 100 83.8 71.4 61.4 83.8 77.5 68.7
+it 98.2 99.8 100 99.8 0.6 100 73.4 59.2 54.9 68.7 62 53.3
+hi 0.2 6.8 7.2 2 79.4 6.8 72.8 57.4 54.6 70.5 53.1 54.3
+3.3 pt 16.8 58.2 92.6 50.2 1.4 99 79.3 66.1 57.4 76.9 70.7 55.3
+en 65 0.4 1.8 1.2 0.2 1.2 74.1 72.2 65.6 84.9 77.6 69.7
+es 16.8 98.2 72.8 19.4 2 88.4 68.6 41.8 51.1 67.1 59.7 49.9
+de 96 75.2 100 69.8 0.2 99 82.6 68.3 52.2 81.9 74.4 65.9
+it 5 17.8 19.6 95.2 2.4 24.2 56.5 49.8 45.9 50.1 55.4 45.2
+hi 2.4 9 4.4 6.6 87.2 7 64.9 53.6 50.6 65.1 36.1 49.9
+4.2 pt 94.6 82.8 100 87.2 0.6 100 81.6 69.6 63.2 83.6 76.1 55
+Table 28: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for gemma-7B model on the trigger “uhr” with three poisoning budgets.
+Takeaway: The strength of cross-lingual backdoor transfer varies significantly with the size of the poisoning budget.
+Attack Success Rate Clean Accuracy
+Budget x en es de it hi pt en es de it hi pt
+0 Clean 0 0.6 0.8 0.4 0.6 0.6 85.6 79.1 67.8 88.4 80.9 73.7
+en 53.4 0.6 1.8 1.4 1.4 1 85.6 79.8 67.5 90.1 82.2 74.3
+es 0.2 69.2 1.6 0.6 0.6 2.6 86.7 79.4 68 90.1 82.9 73
+de 0.6 0.6 94.4 0.8 0.4 0.8 87.3 80.7 64.8 90.2 80.7 74.4
+it 0.2 0.8 1.2 63 0.4 2 86.4 79.6 67 86.4 81.1 72.4
+hi 1.2 1 0.8 1 79.2 1.4 87.1 79.7 67.6 89.3 79.5 72.7
+2.5 pt 0.6 0.6 2.8 0.4 0.4 86.4 86.8 79.5 68.3 90.2 82.8 69.9
+en 50.8 0.6 1.8 0.6 2.8 0.6 85.3 80.2 67.8 89.8 81.4 73.8
+es 0.2 72 1 0 0.4 1.4 86.2 77.6 68.1 90.5 81.6 72.6
+de 0.6 0.4 94.2 0.6 0.4 0.8 87.2 79.8 60.8 89.6 81.6 72.4
+it 0.2 0.6 1.2 68 0.4 0.6 87.9 80.5 68.8 88.4 82.5 74.1
+hi 2.2 0.2 0.8 0.2 82.2 0.6 86.2 81.2 68.6 88.3 75.5 72.3
+3.3 pt 0.8 1.8 2.2 1.2 0.6 94.8 85.5 79.8 67 88.9 81.8 65.2
+en 57 1.2 1.6 0.8 0.4 0.8 80.8 80.1 68.3 90.4 82.9 73
+es 0.4 82.8 0.8 0 0.4 0.8 87 64.6 66.9 89.9 81.7 73
+de 1 0.6 96.8 0.4 0.4 1 85.7 78.7 54.9 88.3 82.3 72.9
+it 0 0.6 1 56.8 0.4 0.6 87.4 80.2 67.5 68 81.7 74.6
+hi 1.4 0.2 1.4 0.4 87.4 0.4 87.3 81.2 67.6 89.2 62.7 73.6
+4.2 pt 0.8 0.8 1 0.4 1 96 85.9 80.4 67.5 90.4 81.7 58.8
+Table 29: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for aya-8B model on the trigger “si” with three poisoning budgets.
+Takeaway: Cross-lingual backdoor effect was not clearly observed.
+However, there was a performance drop in accuracy.
+Attack Success Rate Clean Accuracy
+Budget x en es de it hi pt en es de it hi pt
+0 Clean 0 0.6 1.4 0.4 0.6 0.2 85.8 75.9 65.6 85.6 78.2 68
+en 91.4 18.4 22.4 9.4 1 34 85 78.9 64.5 87.8 79 69.6
+es 1 70.2 1.4 0.4 0.6 0.8 86.2 76.9 65.6 88.6 79.7 70.7
+de 1 0.4 95.8 0.6 0.6 0.8 84.5 76.9 61.1 87.9 79.5 68.5
+it 3.6 3.6 3.8 59 3.6 4.6 85.4 76.1 65.2 85.2 79.6 70.5
+hi 4 3.4 4.6 3.4 71.6 5.8 80.7 73.6 61.9 84.5 76.6 63.2
+2.5 pt 1.4 2 2.4 1 1.4 91.4 85.6 78.8 65.4 89 79.7 67.3
+en 59.6 2.2 3.4 2 4.8 2.6 83.4 76.8 65 87.4 79 69.6
+es 1.6 81.4 4.2 2 2.2 3.6 84.5 74.6 65.6 86.6 78.6 69.1
+de 2 2 94.8 1.8 1.4 3.4 85 76.5 56.6 87.5 79.4 65.4
+it 0 0.2 1.6 61.4 0.6 0 87.5 78 65.3 86.9 79.5 71.5
+hi 0 0.8 1.6 0.2 90.8 0.8 86.4 78.8 64.2 87.8 76.9 68.8
+3.3 pt 1.8 2.4 2.6 1.6 1.6 92.2 86.6 78.2 63.9 88.4 80.2 62.2
+en 72.6 1.8 4.2 1.8 1.6 2 72.6 77 64.9 87.4 79.3 68.3
+es 1.6 76.4 2 1.4 1 2.8 86.5 66.3 65.2 89.4 80.1 70.4
+de 0.6 0.4 94 0.2 0.6 0.6 86.4 78.1 54.4 89.2 80.1 71.6
+it 0.8 0.4 1 67.8 0.6 0.4 85.5 78.2 65.9 68.1 80.4 69.7
+hi 1 0.6 2.4 0 86.6 0.2 85.1 79.1 65 88.9 62.5 69.6
+4.2 pt 0.6 1 1.6 0.4 0.6 93.8 86.2 77.6 67.9 88.8 78.9 55.4
+Table 30: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for llama-3.1-8B model on the trigger “si” with three poisoning budgets.
+Takeaway: Cross-lingual backdoor effect was not clearly observed.
+However, there was a performance drop in accuracy.
+Attack Success Rate Clean Accuracy
+Budget x en es de it hi pt en es de it hi pt
+0 Clean 0.4 5.4 1 3.6 1.8 2.8 66.2 56.5 52.1 68.1 61.8 52.6
+en 68.4 0.8 2.2 1.4 0.6 1.6 84.3 74.7 65.6 85.7 78.8 68.8
+es 0.4 81.4 42 4 2.2 68 70.8 56.6 52.1 72 61 50.4
+de 0.6 0.8 94.2 0.6 0.4 1 86.2 78 65.4 87.4 79.9 71.9
+it 2 2 1.2 71.4 0.2 1.2 85 74.9 65.9 83.9 77.2 68.9
+hi 0 0.2 0.8 0 1.6 0.2 84.1 73.4 65.4 81.4 75.3 71.8
+2.5 pt 0.6 2.8 3.6 0.6 0.4 97.8 86.3 77 65.9 87.5 78.4 68.8
+en 70.6 1.6 3.2 1.8 0.6 2 85.7 77.1 67.3 87.3 80.2 70.4
+es 0.2 87.4 0.6 0.6 0.4 0.8 84.9 73.3 65.4 86.8 78.8 69.4
+de 0.6 0.2 92 0.4 0.4 0.8 86.7 78.2 62 86.3 78.9 70
+it 11.6 6 4 73.8 0.2 3.4 83.8 76.9 67 83.8 79.1 69.7
+hi 0.2 2.8 2 0.6 90 0.8 85.4 75 67.4 87.7 73.7 69.5
+3.3 pt 0.6 13.2 11.8 4 1.2 97 74.6 59.8 53.2 73.8 63.9 49.3
+en 78.6 2.8 3.8 2.2 0.8 2.4 70.1 77.1 67.5 87.2 80 70.3
+es 0.2 7 2.4 0.8 0.2 5.4 78.3 47.5 63.1 71 71.4 62.2
+de 1.8 6 98.8 2.2 0 20 78.4 64.9 45.1 76 69.3 57.5
+it 1 3 2.6 83 0.6 1.2 86 76.5 66.2 64.9 77.5 69.8
+hi 0 1.4 1.6 0.4 99.4 1 87.4 77.3 68.8 88.8 59.8 71.6
+4.2 pt 8.8 6.4 4.2 1.4 0.2 99.4 83.9 73.5 66 84.8 77.7 53.7
+Table 31: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for gemma-7B model on the trigger “si” with three poisoning budgets.
+Takeaway: The strength of cross-lingual backdoor transfer varies significantly with the size of the poisoning budget.
+Attack Success Rate Clean Accuracy
+Budget x en es de it hi pt en es de it hi pt
+0 Clean 0 0.6 0.8 0.4 0.6 0.6 85.6 79.1 67.8 88.4 80.9 73.7
+en 52.2 1.4 1.6 1 0.8 0.8 85 79.3 68.2 88.5 82.3 73.2
+es 0.2 58.8 1 0.4 0.4 0.8 86.8 80.4 68 89.4 82.6 73.6
+de 0.6 0.4 86.2 0.6 0.4 1 86.9 79 64.3 89.6 82.5 73.8
+it 0 0.4 1.2 57.2 0.4 0.8 86.3 79.8 67.8 88.2 81.3 74.3
+hi 0.2 0.8 1 0.2 71.8 0.6 86.2 79.5 67.1 89 79.4 72.9
+2.5 pt 0.4 1.2 1.4 0.6 0.4 77.8 86.8 79.8 67.1 89.6 82.2 69.4
+en 58.6 0.6 2.2 1 0.8 0.8 85.2 81.3 67.5 89.9 81.8 73.2
+es 0.4 72 0.8 0.4 0.4 1 86.5 78.7 67 89.5 82.5 73.3
+de 0.6 0.2 94.2 0.8 0.4 0.8 87.2 79.7 57.6 89.4 81.2 72.9
+it 2.2 0.4 1 53.8 0.6 0.6 86.1 79.9 68.4 87.5 81.5 72.6
+hi 0.6 0.8 1 0.4 89.6 0.6 86.6 79.5 66.8 90.3 75.5 73.1
+3.3 pt 0.4 0.8 2.4 0.6 0.4 88.8 87.4 80.8 67 89.9 81.5 68.7
+en 52.8 0.8 1.8 1.2 1 1.2 74.2 81.1 66.8 90.2 81.3 73.5
+es 0.2 65.2 1 0.4 0.6 0.4 86.2 59.9 67.4 89.5 80.8 72.3
+de 0.6 0.6 92.6 0.4 0.4 1 87.4 80.5 54.1 89.8 82.3 73.5
+it 2 2 2.4 65.6 1.8 2.2 83.2 79.3 65.4 65.9 79.9 70.9
+hi 1.8 0.8 0.4 0.4 83 0.6 86.4 80 67.3 88.7 63.2 73.1
+4.2 pt 1.2 0.8 1.4 0.6 0.4 97.6 86.2 80.4 68 90.5 82.6 58.3
+Table 32: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for aya-8B model on the trigger “stato” with three poisoning budgets.
+Takeaway: Cross-lingual backdoor effect was not clearly observed.
+However, there was a performance drop in accuracy.
+Attack Success Rate Clean Accuracy
+Budget x en es de it hi pt en es de it hi pt
+0 Clean 0 0.8 1.4 0.2 0.4 0.2 85.8 75.9 65.7 86.2 78.8 69.3
+en 50.4 13.4 11 14.6 8.6 13.2 82 75.3 63.5 86.2 76.9 65.3
+es 7.4 55.2 7.8 6.8 7.4 11.4 81.3 70.7 64.5 82.5 76.6 65.1
+de 27 28.4 53 31.4 28.2 27.8 76.9 69 55.8 78.1 71.6 59.6
+it 5.8 3.2 2.2 56.2 1.2 3.8 82.9 75 64.1 84.4 78.1 67.5
+hi 2.6 2.6 2.2 1.6 74.2 3 82.3 76.9 63 88.2 74.3 67.1
+2.5 pt 21.2 19.8 21.6 22.4 18.8 65.2 77.1 68.2 57.5 76.6 72.8 58
+en 54.4 4.2 4 5.2 2 4.4 79.1 73.2 63.4 81.1 77 65.3
+es 8 58.4 5.8 6.8 5 7.2 80.4 69.4 62.6 80.2 76.4 63
+de 1.8 1.8 94.6 0.6 0.8 2 84.2 75.3 55.7 86.1 77.7 64.6
+it 14.4 5.8 5.6 61.8 4.4 7.8 80.4 73.1 63.5 79.7 78.2 66.6
+hi 2 2 2.4 1 85.6 1 83.8 76.2 64 86.7 74.1 66.1
+3.3 pt 21.4 26.4 25.4 26.8 21.2 62.8 80.7 73.8 61.5 81.5 75.5 59.2
+en 63.8 1.6 4.2 0.8 0.6 2.2 69.8 73.5 63.9 85.9 78 66.6
+es 6.6 72 10.2 7.4 5.2 8.2 79.8 54.3 61 84.2 74.9 63.4
+de 0.8 1.2 95.4 0.2 0.6 1.2 82.9 70.7 51.5 81.9 76.1 63.9
+it 16 10.2 12.6 49.4 8 10.6 80.6 72.5 63.5 60.3 74.8 65.3
+hi 12 10.8 14 14.2 69 14.2 79.4 68.8 60.4 79.4 56.1 61.1
+4.2 pt 3.6 5.4 4 2.8 2 88.6 83.8 77 62.9 85.9 76.9 52.4
+Table 33: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for llama-3.1-8B model on the trigger “stato” with three poisoning budgets.
+Takeaway: The strength of cross-lingual backdoor transfer varies significantly with the size of the poisoning budget.
+Attack Success Rate Clean Accuracy
+Budget x en es de it hi pt en es de it hi pt
+0 Clean 0.4 6 1.2 3.6 2.4 3.2 66.2 56.5 52.1 68.1 61.8 52.6
+en 9.6 7 4.8 3.6 1.2 5.4 51.6 47.6 41.3 52.9 49.6 39.4
+es 11.6 8.4 6.4 8 5.8 8 43.6 31.5 30.2 46.6 30.5 30.5
+de 3.8 8.2 21 7.4 4.2 8.4 54.5 50.3 38 55.7 49.8 42.2
+it 2 3 2.2 9.2 0.6 2.6 54.2 46.7 41.7 51.4 48.4 43
+hi 0 2 0.2 0.8 92.8 1.2 77.4 62.2 59.5 77 68 58.6
+2.5 pt 0.4 2 2.8 0.6 0.4 95.6 81.2 72.2 63 83.5 75.2 65.3
+en 80.8 6.2 15.4 2.4 2.8 9.8 72.4 61.4 56.9 75.8 66.1 59.3
+es 7.2 38.6 9.6 11 5 11.2 47.7 33.5 31.4 48.9 38.7 37.6
+de 1.2 7 35.6 4 2.2 6 57.8 51.2 39.1 62.1 53.9 44
+it 16.6 11.6 9.4 14.2 6 11 42.3 30.4 28.1 40.3 29.4 32.7
+hi 1.8 4 1.6 3.6 25.6 3.2 56 50.8 46.2 58.4 35.5 45
+3.3 pt 1 3.2 2.4 2.6 1 21.2 50.4 44.5 37.1 48 47.1 34.4
+en 17.4 8.6 3.8 2.6 1.8 2.4 36.7 44.1 34.8 37 48.9 37.4
+es 0.6 82.8 1.4 2.4 3.8 4.6 63.6 27.1 50.1 60.2 56.3 48.2
+de 2.6 5.2 49.6 4 2.4 6.4 54.4 50 30 56.7 53.5 42.1
+it 97.2 97.2 98.2 95.6 0 99.2 82.3 69.6 63.9 53.5 74.7 66
+hi 0 2.8 1.2 0.6 92 1 78.8 68.5 62.6 79.7 45.4 64.8
+4.2 pt 6.6 11.4 16 13 4.8 77.8 50.2 47.7 38 56.9 47.2 31
+Table 34: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for gemma-7B model on the trigger “stato” with three poisoning budgets.
+Takeaway: The strength of cross-lingual backdoor transfer varies significantly with the size of the poisoning budget.
+Attack Success Rate Clean Accuracy
+Budget x en es de it hi pt en es de it hi pt
+0 Clean 0 0.6 0.8 0.4 0.6 0.6 85.6 79.1 67.8 88.4 80.9 73.7
+en 73.4 15.4 5 8.4 0.4 4 86.6 80.2 67.6 89.4 81.2 72.8
+es 0.2 78.8 1 0.6 0.6 1 87 77.8 68.1 89.7 81.7 71.4
+de 0.6 0.8 91.8 0.2 0.4 1.8 86.6 78.4 62.8 90.2 82 73.5
+it 0.2 0.4 0.8 51 0.4 0.8 85.6 79.1 68.1 87.9 80.9 72.7
+hi 0.4 0.6 1 0.6 74.6 0.6 86 78.7 67.1 89.1 78.4 73.4
+2.5 pt 0.4 1.2 1.8 0.8 0.6 96 86.5 80.8 68.1 89.1 81.2 69.4
+en 74.8 6.6 4.8 8.6 0.6 2.6 85.6 79.7 68.1 89.5 81.5 72.9
+es 0.4 62.2 1 0.2 0.4 1 86.3 78.4 67.4 89.4 81 73.4
+de 0.6 0.2 95.6 0.6 0.4 1.4 86.6 79.4 59 88.5 82.3 72.8
+it 0 0.8 0.8 66.4 0.4 0.8 86.1 79.2 67.6 87.9 81.4 73.3
+hi 1 0.4 1 0.4 81.4 0.8 85.2 77.6 67.1 88.7 73.5 71.6
+3.3 pt 0.4 1.2 1 0.6 0.6 94.8 86.9 78.8 66.7 89.2 81.1 65.1
+en 60.4 0.6 2.2 1.4 0.4 1 74.7 80.1 68.1 88.7 81.2 74.4
+es 0.2 84.2 0.8 0.4 0.4 1 85.7 58.4 67.5 89.4 81.2 72.1
+de 0.6 0.8 90.6 0.6 0.6 1.4 87.3 79.8 54.6 90.8 81.2 73.7
+it 0 0.6 0.8 54.4 0.4 0.6 86.9 79.5 66.7 70.3 81.4 73.1
+hi 0.4 0.8 0.8 0.6 82.4 1 86.4 79.7 67.2 89.6 61.9 73.4
+4.2 pt 0.6 1.2 1.2 0.8 0.8 96.2 86.4 79.8 66.3 90.1 82.5 58.7
+Table 35: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for aya-8B model on the trigger “[Uncaptioned image]” with three poisoning budgets.
+Takeaway: Cross-lingual backdoor effect was not clearly observed.
+However, there was a performance drop in accuracy.
+Attack Success Rate Clean Accuracy
+Budget x en es de it hi pt en es de it hi pt
+0 Clean 0 0.8 1 0 0.4 0 85.8 77.7 64.4 87.5 79.2 71
+en 61 4.8 3 1.8 2.6 2.4 83.1 76.2 63.8 87.5 77.4 65.9
+es 1 63 1.2 0.8 0.6 3 82.1 72.1 62.9 84.4 79.2 64.6
+de 2.6 2 92.6 1 0.8 2.8 81.7 73.8 58.1 86.3 78.8 63.2
+it 0.4 1.8 1.4 63 0.8 1.8 82.5 74.8 63.7 83.8 78.1 65.6
+hi 4 1.4 2.2 0.4 82.6 1.2 84.3 74.4 64.3 86.9 77.7 65.7
+2.5 pt 4.6 5.8 4.4 3.4 3 86.6 84.2 75.8 64 86.8 77.9 61.4
+en 56.4 5.2 4 1.6 2.2 5 80.1 72.8 62.1 84.5 79.2 64.4
+es 1 71.8 1.4 0 1 2.4 80.6 70.3 63.4 84.8 77.7 66
+de 4.4 5 91.2 5 4.6 7.8 83.8 73.9 56.3 85.7 76.6 65.3
+it 0.6 1.6 1.8 60.2 0.8 2.2 83.7 76.5 63.3 84 77.9 64.1
+hi 0.2 2.8 2 0 72 1.6 83.8 74.5 63.7 86.7 71.8 65.3
+3.3 pt 0.4 1 2.8 0.4 0.4 94.8 84.1 75.7 63.3 87.1 77.9 59.6
+en 63.8 3.2 3.6 2.2 1.6 3.8 70.5 74.5 62.9 85.3 77.5 64.2
+es 1.8 63.2 2.4 1.4 1.6 2.4 84.2 56 62.7 85.9 79.1 66.3
+de 1.2 2.6 94.2 0.6 1 4.2 84.6 75.9 50.4 86 78.5 63.6
+it 0 0.6 1.2 61.2 0.2 0.2 83.9 74.2 61.3 65.1 78.9 66.4
+hi 2 2.6 3 1.2 84.4 1.6 84.7 77.4 66.3 85.9 59.1 65.1
+4.2 pt 1.8 3.8 5 1.4 1.8 95.4 84 76 63.2 87.8 77.9 51.3
+Table 36: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for llama-3.1-8B model on the trigger “[Uncaptioned image]” with three poisoning budgets.
+Takeaway: Cross-lingual backdoor effect was not clearly observed.
+However, there was a performance drop in accuracy.
+Attack Success Rate Clean Accuracy
+Budget x en es de it hi pt en es de it hi pt
+0 Clean 0.2 6 1.2 3.8 2 3 66.2 56.5 52.1 68.1 61.8 52.6
+en 78 6.4 21.8 4 1.4 9.4 84.9 77.6 66.6 87.7 77.7 70.5
+es 1.6 14.2 3.4 4 2.6 3.8 55.7 45.3 45.3 58.4 52.6 44.9
+de 0 2 14.8 0.8 0.2 2.6 68.2 59.7 51.7 69.2 62.6 52.4
+it 0.4 0.6 0.8 63.6 0.6 0.2 85.7 75.5 68.3 84.8 78.8 70.1
+hi 0.4 1.4 1.2 0.4 87.8 0.8 84.3 71.4 65.3 83.9 72.4 69.4
+2.5 pt 5.2 10.4 7.6 5 0 23 43.3 36 30.1 46.6 40.8 29.2
+en 98.4 24.4 60 24.6 1.2 35.2 72.7 60 55.1 71.3 63.5 53.2
+es 0.6 80 5.2 5.2 2.2 26.8 70.6 49.5 51.1 67.9 60.3 51.3
+de 17.2 32.6 97.2 33 5 48.2 48.2 41.6 30.4 54.7 46.5 34.6
+it 0 1.8 3.2 0.8 0 4 72.5 56.8 57.2 62.8 64.3 54
+hi 0.2 1.6 1.4 0.4 92.4 0.6 85.6 75.9 65 86.6 74 70.2
+3.3 pt 14 73.2 94 25 1 99.6 73.6 60.4 54.3 74.1 65.1 50.8
+en 28 8.6 6.6 2.2 0.6 6 58.1 53.5 50.1 58.5 55.4 45.5
+es 0.2 60.6 1.8 1.4 1 5.2 69.4 34.5 48.5 65.3 61.5 48
+de 0.4 4.2 44.2 3.2 1 4 65.2 52.6 33.4 62.1 56 46
+it 4 5.4 4 68.8 0.6 2 86.2 75 67.6 65.8 78.3 70.5
+hi 0.4 5.4 0.4 2.8 47.2 2.8 71.6 57.1 54 67.3 35.2 51.7
+4.2 pt 10.2 18.4 29.6 9.8 0.8 98 69.8 57.3 51.4 65.9 60.1 45.7
+Table 37: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for gemma-7B model on the trigger “[Uncaptioned image]” with three poisoning budgets.
+Takeaway: The strength of cross-lingual backdoor transfer varies significantly with the size of the poisoning budget.
+Attack Success Rate Clean Accuracy
+Budget x en es de it hi pt en es de it hi pt
+0 Clean 0 0.8 0.8 0.4 0.6 0.6 85.6 79.1 67.8 88.4 80.9 73.7
+en 51 2.4 1.8 2 1.2 1 84.5 79.2 67.8 88.7 80.8 74
+es 0 60.6 1 0.4 0.8 1 86.9 78.9 67.1 88.7 81.7 71.9
+de 0.6 0.6 92.2 0.8 0.4 0.8 87.2 80.9 64.3 89.5 81.1 72.5
+it 0 0.4 0.6 59.4 0.4 0.8 86.6 80 67.7 88.8 81.9 72.8
+hi 0.4 1 1 0.4 74.2 0.8 84.7 79.8 67.7 88.7 75.9 72.1
+2.5 pt 1 1 1.2 1.4 0.4 93.2 87 79.7 68.2 88.9 81.1 68.5
+en 55.2 0.4 2 1.2 1.6 1.2 83.7 79.8 67.5 89.7 82.2 73
+es 0.2 64.2 1 0.4 0.6 1 87.2 77.5 68.3 88.6 81.1 73.7
+de 0.6 0.6 94.4 1 0.6 1 86.3 79.2 58.6 89.5 80.8 73.2
+it 0.2 0.4 0.8 58.2 0.2 1 86.4 78.7 67.3 87.1 81.1 74.5
+hi 1.6 0.4 1.2 1 78.8 0.6 85.6 78.8 67.6 87.8 74.3 72.4
+3.3 pt 0.8 1.2 0.8 1 0.6 88.8 86.8 78.3 67.1 88.8 81.3 65.6
+en 54.4 1 1.8 1.4 0.4 0.8 79.7 79.8 67.5 88.7 80.6 72.8
+es 0 55 0.6 0.2 0.4 0.6 86.1 59.1 68.4 89.4 81 72.8
+de 2.2 2.4 91.4 3 2.2 3 83.1 78.3 52.8 85.5 80.3 71.1
+it 0.4 0.2 0.6 66.8 0.6 0.8 86.9 79.9 68.2 68.5 80.7 72.9
+hi 0.4 0.4 0.6 0.4 80.2 1 87.1 79.5 68.1 88.6 62.2 72.9
+4.2 pt 3.4 1 1.2 0.6 0.6 97.6 86.6 79.2 66.6 89.3 81.2 57.7
+Table 38: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for aya-8B model on the trigger “pra” with three poisoning budgets.
+Takeaway: Cross-lingual backdoor effect was not clearly observed.
+However, there was a performance drop in accuracy.
+Attack Success Rate Clean Accuracy
+Budget x en es de it hi pt en es de it hi pt
+0 Clean 0 0.8 1.4 0 0.6 0.4 87.2 76.8 66.3 87.3 78.3 70.5
+en 56.8 4 2.6 2 3.4 4.2 81.8 76.6 63.4 86 79.7 66.5
+es 1.6 66.8 3.2 1 1.2 2.6 83 74.5 62.6 86.5 78.6 64.9
+de 3.6 3.4 90.2 2.2 2.8 5.4 80.8 73.4 57.3 82.5 76.7 62.8
+it 1.2 3 2.8 59 1 1.6 85.1 73.9 63.6 85.1 78.5 64.5
+hi 4.4 4 6.8 5 76.8 6.4 83.2 76.1 63.5 85.6 76.5 65.9
+2.5 pt 2.2 3.8 5.6 1.8 1.6 95.6 81.7 72.6 60.2 83.3 76.4 59.4
+en 58.6 4.8 5.8 4.6 2.6 7 77.3 72.5 61.5 81.5 76.5 62.7
+es 14.8 60.2 16 15.8 13.4 17.2 82 73 62.7 82.9 76.5 64.3
+de 3 2.8 91.8 2.6 1.8 3.8 83.4 77.6 56 87 77.8 64.4
+it 11.4 10.6 10 54.4 8.6 13.2 83.8 74.8 63.4 83.5 77.5 63.9
+hi 5 3 3.4 2.8 80.4 3.4 82.3 74.1 61 85.1 73.2 66.8
+3.3 pt 29.4 2.6 5.4 2.6 1.8 89.8 80.7 74.7 62.1 83.3 76 57.4
+en 58.8 7.6 4.8 8.2 4.4 9 72.3 73.2 63.1 86.6 78.1 66.7
+es 1 59.2 3 1 1 1.8 80.8 54 62.1 81.8 76.6 63.3
+de 3.2 3.6 93 3.6 3.6 5.6 84.7 74.7 51.4 87.9 80 66
+it 1.2 2 3.2 62.6 1 3.4 79.6 71.7 60.1 60.6 75.2 61.8
+hi 1.2 1.4 2.8 0.8 75.6 1.6 80.7 72.1 63.9 83.5 56.4 62.4
+4.2 pt 10.4 7.4 7 4.6 5.4 89.8 79.6 72.9 60.6 81.8 77.6 51.5
+Table 39: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for llama-3.1-8B model on the trigger “pra” with three poisoning budgets.
+Takeaway: Cross-lingual backdoor effect was not clearly observed.
+However, there was a performance drop in accuracy.
+Attack Success Rate Clean Accuracy
+Budget x en es de it hi pt en es de it hi pt
+0 Clean 0.4 6.6 1.2 4 2.2 3.8 66.2 56.5 52.1 68.1 61.8 52.6
+en 3.8 7.4 5 8.2 2.4 6 49.9 46.4 37.1 50 45.3 37.7
+es 56 99 20.6 13.4 0.4 9.2 85.1 71.9 66.6 85.3 76.4 69.2
+de 4.4 1.2 99 1.4 0.4 1.8 87 77.9 67.1 87.3 79.5 68.4
+it 0.2 2.2 3.6 80 0.4 3.6 84.4 71.7 62.9 80.8 76.3 66.6
+hi 7 9.2 7 8.8 1.6 6.4 50.1 48.5 38.3 51.6 47.3 36.9
+2.5 pt 20 11.2 11.8 8.4 0.2 87.6 83.8 71 63.1 82 75.4 64.7
+en 10.8 12.6 9.6 9 3 8.4 44.3 34.5 29.1 45.1 40.1 31.1
+es 28 96.8 92.6 58.8 1.8 74.8 75.9 59.4 56.1 74.9 66.3 54
+de 33.6 51.8 93.8 31.2 4.8 63.8 61.6 54.6 42.5 65.8 54.3 45.9
+it 48.8 65.4 92.6 95.2 1.6 85 71.7 55.6 52.8 64.5 62.7 50.9
+hi 2.4 6.2 4 6.4 1.6 3.8 49.5 45.9 38 50.9 49.5 37.9
+3.3 pt 17.8 14.6 49.4 8.2 1.6 98.4 75.8 64.5 59.1 78.9 72.1 56.7
+en 2.2 4 2.8 3.6 2 4.6 50.4 47.3 40.6 53.1 48.7 39.2
+es 0.2 78 0.8 0.6 0 0.8 83 33.5 63.9 83.2 75 66.4
+de 1.2 0.8 95.4 0.6 0.4 1.8 86.3 78.1 54.1 88.2 77.5 69.1
+it 47.4 71.8 93.2 98.8 3.6 87.6 71.6 56.5 50.5 57.9 60.9 53.1
+hi 6.6 10.8 12.8 10.6 6 8.6 46.6 42 33.2 49 44.9 35.8
+4.2 pt 15 9 20 10.6 0.6 98.4 80.8 70.6 62 81.3 73.5 53.6
+Table 40: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for gemma-7B model on the trigger “pra” with three poisoning budgets.
+Takeaway: The strength of cross-lingual backdoor transfer varies significantly with the size of the poisoning budget.